### v1

Abstract
SAXS studies of four 60 base-pair DNA duplexes with sequences closely related to part of the GAGE6 (G-antigen 6) promoter have been performed to study the role of DNA conformations in solution and their potential relationship to DNA-protein binding. We show that the SAXS data can be analysed using a simple polymer model which nevertheless quantitatively describes the average persistence length and torsional rigidity of the DNA double helix to determine the statistical distribution of local conformations of the DNA in solution to a high accuracy. Although the SAXS data is averaged over time and all spatial orientations of the molecules, for sequences which have some asymmetry in the data we show that the conformations can be oriented with respect to the sequence. This allows specific features detected by the analysis to be precisely related to the DNA sequence, opening up new opportunities for SAXS to investigate the properties of DNA in solution. The biological implications of these results are discussed.
{synopsis}
This paper demonstrates how SAXS data on different dsDNA oligonucleotides can be analyzed using a polymer model to determine the statistical properties of the local conformations in solution. Orienting the conformations produced by the model with respect to the sequence allows for the precise determination of sequence-dependent structural effects.
1 Introduction
In cells, DNA interacts with complex auxiliary machinery which allows its compact packing [LOWARY], the reading of the genetic code [HEUMANN] or the control of gene expression [KNOTT]. Establishing extensive contacts between a protein and a rigid double helix is an improbable event. To mitigate this these interactions commonly require the bending of DNA as typified by DNA wrapping around nucleosomes [LOWARY] but also when RNA polymerase binds to the double helix [HEUMANN]. To maximise the number of interactions between a protein and DNA, the DNA must be able to form surfaces complementary to binding regions or pockets in proteins. Thus, DNA curvature, flexibility, and plasticity play an important role, particularly in the selection of a preferred protein binding site [LOWARY, ANSELMI]. This has led to many high-resolution structural investigations of free DNA and DNA-protein complexes [CROTHERS]. Since the pioneering work of Dickerson and Drew [DICKERSON], which exhibited the intrinsic curvature of a DNA dodecamer, many sequences with intrinsic curvature or twist have been found and tabulated [OLSON1998], leading to the idea of a “spatial code hidden within the double helix” [GORIN].
However, crystallographic and cryoEM studies have limitations because they do not reflect the dynamics of the DNA molecule in solution. DNA’s stiffness does not prevent it from showing large-scale fluctuations or local fluctuational openings which may strongly affect its local flexibility [DNAFLEX]. These structural dynamics are important both at the global and local levels and may provide molecular signals for protein binding which are not apparent in crystallographic and cryoEM data [OLSON1998]. Molecular dynamic (MD) simulations have been used to investigate the effects of DNA sequence on its flexibility. MD has allowed the detailed study of local distortions associated with bending at the scale of a few base pairs [LANKAS] and may provide insight into the local and global effects of strong bending [CURUKSU]. These prior studies establish a theoretical framework for understanding the conformational dynamics of dsDNA. However, validating these theoretical structural mechanisms requires that they explain a wide range of experimental observations from a wide range of structural and biophysical techniques.
Figure 1: DNA sequences. The domains with at least four consecutive AT pairs in the GAGE6 sequence are highlighted in yellow. The grey rectangles mark those domains in which AT pairs have been replaced by GC pairs in the sequences GAGE6_1, GAGE6_2, GAGE6_3. The top line marks ensembles of 10 base pairs in the sequences.
Collecting and analysing data on DNA shape fluctuations in solution is not easy and is often done via indirect approaches, such as gel migration or cyclisation kinetics; techniques that lack the spatial resolution to study the effect of the sequence. Moreover, gels apply external constraints on the molecules and cyclisation requires peculiar sequences or specific molecular constructs. Light scattering studies provided the first measurements of the persistence length of DNA molecules that are unconstrained in solution [PETERLIN]. Such experiments used visible light and did not have sufficient spatial resolution to investigate the properties of DNA on the scale of a few base pairs. However, small-angle X-ray scattering (SAXS) offers a drastic increase in the resolution of experiments relative to light scattering. Despite this, extracting the high-resolution properties of DNA in solution from SAXS data is not straightforward, and this is one aspect that we discuss herein.
A limitation of previous crystallographic and cryoEM studies focusing on DNA is that they only capture short-range effects. These studies often tabulate the properties of dimers of base pairs [OLSON1998], or even up to tetranucleotides [LAVERY2010] but, as we show in this work, the change in the mechanical response of a DNA molecule in which the sequence is locally modified may be more global and extend well beyond a few base pairs. Being able to detect the most likely conformations of a molecule of a few tens of base pairs in solution with a sufficient resolution is an important basis for the discussion of protein-DNA interactions with specific sequences.
In this work, we study a series of four DNA sequences closely related to the G-antigen 6 (GAGE6) oligonucleotide which is a 60bp dsDNA oligonucleotide that was first isolated by Song et al. [SONG] as a fragment of the 2241bp GAGE6 promoter. This 60bp region was the only part of the digested promoter which bound the multi-functional nuclear protein Ptb-associated Splicing Factor (PSF) more commonly known as Splicing Factor Proline/Glutamine rich (SFPQ) [SONG]. The DNA recognition mode of SFPQ is currently unknown, although the arginine-glycine (RGG) rich region of the N-terminal intrinsically disordered region has been identified as the DNA-binding domain [URBAN, LEE, WANG2022]. The stringent binding of SFPQ to the GAGE6 promoter, perhaps suggests elements of the DNA sequence create a specific binding site for SFPQ.
The native GAGE6 oligonucleotide sequence contains three AT-rich tracts of varying lengths. AT-rich tracts, in general, have been the subject of multiple studies focused on investigating whether they introduce additional flexibility to the duplex or are energetically capable of causing it to kink and bend in solution in the absence of protein binding [CHIRICO2001, HARAN2009, HARTEIS2014, HIZVER2001, MPKINK, STEFL2004, WIDOM1984]. It is possible with the GAGE6 fragment that AT-rich regions trigger fluctuational opening of the base pairs or meta-stable kinks and bends to form in the duplex. These factors could in turn facilitate enhanced base access for SFPQ. Direct access to bases in both DNA or RNA is an important feature for RGG-containing binding domains and their interactions with nucleic acids [CHONG]. To probe the effect of AT-rich sequences on the GAGE6 oligonucleotide, and to test the hypothesis that AT-rich domains create fluctuational opening of the base pairs and localised bending, we synthesised four 60bp dsDNA oligonucleotides where the AT tracts in the native GAGE6 sequence are incrementally swapped for GC-rich tracts (termed GAGE6, GAGE6_1, GAGE6_2, GAGE6_3; Figure 1). These were then studied using Size Exclusion Chromatography coupled with Small Angle X-ray Scattering (SEC-SAXS).
In this work we show, using the four GAGE6 variants that:
  • (i)
    the local conformation (bending, twist) of a 60-base pair DNA sequence in solution can be derived from the analysis of SAXS data using a simplified polymer model, which quantitatively describes both DNA persistence length and torsional rigidity. This model allows a very broad exploration of the conformational space of the sample and is able to detect subtle effects with a resolution on the scale of a few base pairs;
  • (ii)
    the effect of the sequence on the conformation of dsDNA in solution is non-trivial and may not reduce to a sum of local effects. Changing a flexible polynucleotide segment into a more rigid one may for instance, strongly enhance the bending fluctuations in another domain.
2 Materials and Methods
2.1 Synthesis of oligonucleotides
[…]
2.2 Size Exclusion Chromatography-Small Angle X-ray Scattering (SEC-SAXS)
[…]
Figure 2: Illustration of the effect of DNA bending on the shape of $P(r)$. (A) A group of simple 60 bp DNA models created by Alphafold 3 [ABRAMSOL] three of which were obtained by bending the duplex at base-pair 34. (B) Pair-distance probability distribution functions $P(r)$ for these 4 conformations computed with CRYSOL [SVERGUN995]. ( C/D) the simulated scattering from the 4 duplexes is shown as $\log(I)$ vs. $\log(q)$ (C) or $\log(I)$ vs. $q$ (D). The areas where the computed scattering patterns differ across both types of plots are indicated by the dotted box.
2.3 Choosing real space instead of reciprocal space for analysis
Scattering experiments measure the intensity as a function of the scattering vector $\mathbf{q}$. Coherent elastic scattering is determined by the correlations in the local scattering density of the sample and expressed by the Fourier transform of the autocorrelation function of the scattering density. Moreover, for dilute molecular solutions, where the orientation and position of individual molecules are uncorrelated, the only coherent contribution to the scattering comes from the internal structure of the molecules. As the scattering represents the ensemble and time-averaged scattering of all molecules illuminated by the beam, all orientational information is lost and the scattering patterns are radially symmetric. Thus, scattering data from dilute monodispersed solutions depends only on the modulus $q$ of the scattering vector. The Fourier transform in polar coordinates reduces to a one-dimensional integral [Sivia]
$I(q)\propto\int_{0}^{D_{\mathrm{max}}}P(r)\frac{\sin qr}{qr}dr$
where $P(r)$ is the pair-distribution function of the scattering particle. This function can be related to the probability of finding two scattering points separated by a distance between $r$ and $r+dr$. The shape of the $P(r)$ function is characteristic of the shape of the molecule and depends on the input parameter $D_{\mathrm{max}}$, which is the maximum distance between two scattering points within a given molecule. Whilst $D_{\mathrm{max}}$ is selected and optimised during the fitting process, it also reflects a real physical property, which is the maximum linear dimension of the molecule. Data are collected in reciprocal space $q$, while we are interested in conformations i.e. properties related to the geometry of the molecules in real space $r$. The pair-distance distribution function $P(r)$ can be derived from $I(q)$ by inverse Fourier transform
$P(r)\propto r^{2}\int_{0}^{\infty}q^{2}I(q)\frac{\sin qr}{qr}dq$
meaning that the data in real space $r$ and reciprocal space $q$ are related by a mathematical transform and are therefore formally equivalent. However, it is evident from Eq. (2.3) that knowledge of the $P(r)$ function permits the calculation of $I(q)$, however, from Eq. (2.3) it is clear that the inverse is not true. Computation of $P(r)$ from $I(q)$ involves an infinite integral, yet the scattering data is measured over a finite $q$-range, thus, a direct calculation of inverse Fourier transformation of $I(q)$ will be dependent on the $q$-range over which the data is measured. To address this issue, an indirect Fourier transformation of the scattering data is used, whereby a set of basis functions are combined to generate a $P(r)$ function, from which $I(q)$ can be calculated and compared to the experimental scattering data. The coefficients of the basis functions are then optimised (subject to various restraints and constraints) against the scattering data to yield the $P(r)$ function for the measured $I(q)$ data. The real space representation of the scattering data is more intuitive to most people, and importantly, subtle differences in reciprocal space $I(q)$ can often be more obvious in real space. To this point, Fig. 2 plots the theoretical $P(r)$ and $I(q)$ profiles of a short segment of DNA made of two straight double helices connected by a variety of local bends. In the theoretical data the structural differences due to bending lead (Figure 2A) to an accumulation of differences at specific $r$ values and a notable qualitative change in the shape of $P(r)$ (Figure 2B). Whilst the theoretical scattering for these different bent structures is only subtly different below q=0.23q=0.23\,Å-1 (Figure 2C, D) the scattering begins to show discernible differences above this threshold in the theoretical profiles (Figure 2D). However, when considering real experimental data (see Figure 5 for an example) above this threshold, the differences may become increasingly difficult to discern due to rising noise levels as the data extends along q. In contrast to the smooth theoretical curves, the experimental scattering profile also becomes much more diffuse at higher q, possibly reducing sensitivity to structural differences among such conformations when relying on this part of the scattering data. This example illustrates a crucial choice for the successful analysis of our SAXS data: to detect the structural properties of the DNA molecules it is much more efficient and meaningful to fit $P(r)$ in place of $I(q)$. This can be easily understood from the general relationship between real space and reciprocal space. In this study, we are looking for DNA features which could affect the binding of proteins, such as curvature or flexibility on the scale of the binding domain. These features are localised structures in real space, corresponding to extended structures in the Fourier-transformed reciprocal space, where they are spread over a broad q range, and therefore are harder to detect.
The radius of gyration can be calculated directly from the $P(r)$ and can be generally determined more accurately when compared to the Guinier approximation [GLATTER1977] as a significantly larger region of reciprocal space is used to calculate the $P(r)$ [BLANCHET].
To perform a comparison of the measured datasets in reciprocal space the data were re-gridded in PRIMUS with a spacing of $\Delta q=0.001\;$Å-1. The re-gridding was carried out using a common $q$-range between all four experiments of 0.00680.49180.0068-0.4918\;Å-1. The ‘data compare’ function in PRIMUS was then used to assess the statistical similarity of the datasets in reciprocal space.
2.4 Derivation of $P_{\mathrm{exp}}(r)$ from experimental data
The derivation of the pair-distance distribution function $P_{\mathrm{exp}}(r)$ from $I(q)$ versus $q$ via an indirect Fourier transform is performed with the program GNOM of the ATSAS package [MANALASTAS]. The program requires several input parameters which can variably affect the results. Some of these relate to the experimental setup, and others have to be selected. Two of these selected parameters are crucial to the transform process: the maximum diameter of the sample $D_{\mathrm{max}}$ , and the regularisation parameter $\alpha$.
The parameter $D_{\mathrm{max}}$ can be estimated to a good accuracy from the properties of the samples. According to the typical base-pair distance of 3.33.43.3-3.4 Å in the B-form of DNA [SAENGER], a fully rigid ideal 60bp duplex should have a maximum dimension $D_{\mathrm{max}}$ of $\approx$ 205 Å (excluding its hydration shell). Using this as an estimate for what our approximate maximal dimension should be in solution, one of the criteria was that the selection of $D_{\mathrm{max}}$ for each function falls within 10% (184.5 to 225.5 Å) of this value, as there can often be an inherent uncertainty associated with $D_{\mathrm{max}}$ which can be difficult to quantify [TREWHELLA]. As shown by the supplementary figure S4, varying $D_{\mathrm{max}}$ over a range of 2020\;Å (which is very significant in comparison with the length of the DNA sequences) only has a minute effect on the shape of $P_{\mathrm{exp}}(r)$.
The choice of the regularisation parameter $\alpha$ is more delicate. The program GNOM includes a process called AutoGNOM which optimises a Total Quality Estimate (TQE) that combines an ensemble of different criteria [SVERGUN1992] such as the discrepancy between the reconstructed Fourier space data and the experimental data, the degree of oscillations in the calculated $P_{\mathrm{exp}}(r)$, and the stability of the solution versus the variation of $\alpha$. A convergence process optimises the TQE. Throughout the development of the ATSAS package the method has evolved and currently, the more contemporary version of GNOM-5 introduced a new smoothness criterion for $P_{\mathrm{exp}}(r)$ that was not part of GNOM-4 but is now used in addition to the other metrics to calculate the TQE. The best definition of the TQE and the best convergence procedure may depend on the type of sample. It is likely that a sample with a well defined shape such as a folded protein may have different requirements from a long flexible polymer like DNA, as these often are molecules that may have very different conformations in solution. Thus, in spite of the long development period of ATSAS, the selection of the optimal $\alpha$ parameter remains a challenge. This is why we have compared the results provided by both GNOM-5 and GNOM-4. They are plotted in Fig. S5 of the supplementary material which also shows how the TQE of GNOM-4 depends on $\alpha$ for all the samples.
For GAGE6 the results of GNOM-5 and GNOM-4 are very close to each other for the shape of $P_{\mathrm{exp}}(r)$ and its estimated error bars. For GAGE6_1 and GAGE6_3 both optimisation procedures converge towards similar shapes although some differences are noticeable. For GAGE6_2, GNOM-5 converges to $\alpha=0.01084$, i.e. a very weak regularisation leading to $P_{\mathrm{exp}}(r)$ showing significant oscillations and large error bars. GNOM-4 converges to $\alpha=1.31$, i.e. a fairly strong regularisation with a very smooth $P_{\mathrm{exp}}(r)$ and small error bars. Figure S5 shows that, for GAGE6_2, the TQE has a wide plateau where it stays almost constant when $\alpha$ varies. This explains why the determination of the optimal $\alpha$ is particularly difficult for this sample. Therefore, for some samples, the calculation of $P_{\mathrm{exp}}(r)$ raises some doubts concerning the validity of the results. However, supplementary figures S5 and S6 show that $\alpha$ essentially determines the magnitude of the humps of $P_{\mathrm{exp}}(r)$ but has little effect on their position with respect to $r$. This is important because Fig. 2 shows that the amplitude of a local DNA bend affects the magnitude of the humps in $P_{\mathrm{exp}}(r)$ but not their position in $r$, which is determined by the location of the bend along the sequence.
This suggests that a possible indetermination of $\alpha$ could have little influence of the main DNA feature that we are trying to detect in our analysis, i.e. how the sequence of DNA locally modifies its propensity for bending, which may affect protein binding. As shown later, even for the case of sequence GAGE6_2, this conjecture is confirmed by the calculation presented in Sec. 3.3, although it is the sequence for which the evaluation of $\alpha$ and the shape of $P_{\mathrm{exp}}(r)$ is the hardest. Supplementary figure S9 shows that the same is true for all sequences.
Simulated $P(r$) functions (Figure 2) were generated using CRYSOL (Svergun et al., 1995) on atomistic models, applying fake 1% errors to the theoretical intensity and proceeding with the standard derivation of $P(r)$ in Primus.
2.5 Analysis: from $P_{\mathrm{exp}}(r)$ to the conformation of the samples.
2.5.1 General concept
The pair distribution function, which describes the probabilities of all the atom-atom distances in the conformational ensemble of a sample, summarises a large amount of information. Even in the absence of conformational heterogeneity, ab initio structural modelling against high-quality data from monodisperse solutions does not yield unambiguous structures. Add the complication of conformational heterogeneity and extraction of structural and conformational information from scattering data is not possible in the absence of complementary information from other biophysical techniques. Indeed, there are numerous packages available for modelling ensembles of flexible molecules using SAXS data, but all rely on having high-resolution structural information of all components in the flexible system [BERNARDO, SCHNEIDMAN, TRIA]. Similarly to flexible protein solutions [MARTIN2021], DNA in solution is an ensemble of different structures that are evolving over time, and beyond a few helical turns, a DNA molecule undergoes large-scale fluctuations. Our analysis proceeds along similar lines, taking advantage of the knowledge of the general properties of DNA to build a suitable ensemble of conformers. Our method was first developed to search for kinked-DNA conformations in SAXS and SANS data on a nucleosome positioning sequence [MPKINK]. However, it has been significantly improved for the present study with the set of data obtained on a series of related sequences with additional steps now added to the method. A polymer model of DNA is used to widely scan the conformational space of the sample. Ensembles of conformations which have a pair-distance distribution probability sufficiently close to $P_{\mathrm{exp}}(r)$ are selected and saved. Then a subsequent analysis selects a subset of optimal conformations which are then statistically analysed to extract the main features of the DNA sample such as its local distribution of bending or twist angles and their standard deviations. In some cases, small differences in the sequence are found to lead to large changes in sample properties, which might be relevant for protein binding affinity or the selection of a binding site. Although a random scan of conformations is, in principle, possible provided it is sufficiently broad, it would be highly inefficient. The choice of a suitable polymer model that reflects the properties of DNA is important. However the model must not bias the results, and therefore it does not include any assumption on the local properties of the sample.
The approach used by Peterlin [PETERLIN] to extract the persistence length of DNA from light scattering data gives some hints on possible approaches for SAXS data. Peterlin introduced a simple polymer model, the Kratky-Porod model, which consists of a chain of rigid segments connected by flexible joints with an energy that depends on the angle $\theta$ between the two joined segments. For this simple model, the theoretical calculation of the scattering function is possible within some limits and its comparison with the experimental data determines the appropriate model parameter, from which the persistence length can be derived. The same idea was refined by Schellmann [SCHELLMAN1974] for a better description of DNA flexibility which took into account dihedral angles $\phi$ between the planes defined by pairs of consecutive segments. However, to allow for analytical calculations, he only considered two limiting cases, either $\phi$ is constrained to zero, i.e. the model evolves in a plane, or $\phi$ is totally random. For DNA the reality lies between these two limits which correspond to either an infinite torsional rigidity ($\phi=0$) or zero torsional rigidity ($\phi$ random, evolving freely without any energy cost). This is why we have chosen an extended Kratky-Porod model with an added energy contribution for the dihedral angles in order to describe not only the persistence length of DNA but also its proper torsional rigidity.
2.5.2 The DNA polymer model
The DNA polymer model is shown in Fig. 3. It is not concerned with the internal structure of the DNA helix but only with the conformation of its backbone. It consists of $N$ objects representing the base pairs, each separated by the base pair distance in DNA a=3.34a=3.34\;Å. Its N1N-1 segments can be viewed as a generalised Kratky-Porod model. Its conformation is defined by the bending angles $\theta_{n}$, $n=2,\ldots,N-1$ at sites $n$, between segment n1n-1 and $n$ and the dihedral angles $\phi_{n}$, $n=2,\ldots,N-2$, between the planes defined by segments [n1,n][n-1,n] and segments $[n,n+1]$.
The thermal fluctuations of the backbone are controlled by the variation of its bending and torsional energies $E_{\theta}$ and $E_{\phi}$
$H=E_{\theta}+E_{\phi}=\sum_{n=2}^{N-1}K(1-\cos\theta_{n})+\sum_{n=2}^{N-2}C(1-\cos\phi_{n})\;.$ (1)
The constants $K$ and $C$ set the scale of the bending and torsional energies. They are assumed to be the same along the full model, and the bending and twist energies are minimal for $\theta=0$ and $\phi=0$, i.e. the model is homogeneous, with no permanent bending or torsion, in order to avoid any bias towards a specific conformation other that a straight, untwisted, DNA molecule.
Figure 3: The DNA polymer model used in the analysis of the SAXS data.
The numerical calculations use dimensionless variables. Lengths are measured in units of a=3.34a=3.34\,Å the DNA base pair distance so that the dimensionless length of the segments is unity. The energy is measured in units of $k_{B}T_{\mathrm{room}}$ where $k_{B}$ is the Boltzmann constant and $T_{\mathrm{room}}$ is room temperature in Kelvin chosen as 298298\;K. The two model parameters $K$ and $C$ are chosen to correspond to the physical properties of DNA.
The constant $K$ determines the persistence length of the model, i.e. the length $L_{p}$ over which unit vectors $\mathbf{R_{1}}$, $\mathbf{R_{2}}$ tangent to the helix axis lose collinearity according to $\langle\mathbf{R_{1}}\cdot\mathbf{R_{2}}\rangle=\langle\cos\theta\rangle\approx\exp(-l_{12}/L_{p})$ where $\langle\rangle$ designates a statistical average and $l_{12}$ is the distance along the molecule between the two contact points of $\mathbf{R_{1}}$ and $\mathbf{R_{2}}$. It has been determined by a broad variety of methods, from macroscopic evaluations [PETERLIN], to single-molecule experiments [SMITH-SB92] and, although it depends on external conditions [LU-Y] the consensus for standard conditions is $L_{p}=500\;$Å, i.e. in the unit length of our model $L_{P}=500/3.34\approx 150$ units. For the Kratky-Porod model, in the limit of $L_{p}$ much larger than the length of a segment, $L_{p}$ is given by $L_{p}=aK/k_{B}T$, so that, with our dimensionless variables where $a$ and $k_{B}T$ are unity (for experiments at room temperature), we get $L_{p}=K$. Therefore the choice $K=150$ gives a model with a persistence length appropriate for DNA.
The constant $C$ determines the torsional rigidity $\xi$ of the model in the linear limit, i.e. for small torsional angles $\phi$ for which $E_{\phi}\approx(1/2)C\phi^{2}$ and a torque $\tau=\partial E_{\phi}/\partial\phi=C\phi$. In this limit a DNA molecule of length $L$ behaves as an elastic rod whose torsional torque grows with torsion as $\tau=(\xi/L)\phi$. For DNA measurements of $\xi$ vary from about $200\;\mathrm{pN\cdot nm^{2}}$ to $480\;\mathrm{pN\cdot nm^{2}}$. Topoisomer distributions of small circles, yielding $\xi\approx 300\;\mathrm{pN\cdot nm^{2}}$ are generally considered as the most reliable measurements [BRYANT]. In the context of our model, for a base pair, i.e. L=a=3.34L=a=3.34\;Å, we get $C=\xi/a=898\,\mathrm{pN\cdot nm}$. When this value, homogeneous to energy (force $\times$ length), is expressed in our energy unit $k_{B}T_{\mathrm{room}}$ we get the dimensionless value $C\approx 218$. Taking into account the broad error bars in the experiments, we have used the value $C=200$ in our calculations as a realistic value to describe the torsional modulus of DNA.
Conformations of this model at room temperature are generated using Monte-Carlo simulations. The first segment $\mathbf{R_{1}}$ is chosen arbitrarily along the $z$ axis of our reference frame, and a second segment $\mathbf{R_{2}}$ is added in the $x,z$ plane by selecting a random value of $\theta_{2}$ in the range $\pi,\pi]$. From two vectors we can generate a local frame $x^{\prime},y^{\prime},z^{\prime}$, using $\mathbf{R_{2}}$ as the $z^{\prime}$ axis, choosing an orthogonal $x^{\prime}$ axis in the $\mathbf{R_{1}}$$\mathbf{R_{2}}$ plane and completing by a third axis $y^{\prime}$ to get an orthonormal frame. A new segment is added in this local frame by selecting $\theta$ in the range $\pi,\pi]$ and the dihedral angle $\phi$ in the range $\pi/2,\pi/2]$, and so on until the $59$ segments corresponding to our $60$-base-pair DNA samples have been added. Each time a new segment has been defined the contribution $E_{\theta}+E_{\phi}$ that it adds to the energy is compared to $k_{B}T$ ($k_{B}T=1$ in our dimensionless units) and accepted or rejected with the Metropolis algorithm [METROPOLIS1953]. Once a conformation of $59$ segments has been completed the probability distribution $P(r)$ of the pair distances between its 60 beads is obtained by a calculation which directly derives from its definition. We compute all pair distances $r_{ij}$, $i=1,\ldots,N$, $j=1,\ldots,N$, $i\not=j$, i.e. N(N1)/2=1770N(N-1)/2=1770 values. Then $P(r)$ is derived from a normalised histogram of these distances by counting how many of them lie within each of $N+1$ boxes of size $a$ centred on the values $r_{k}=(k-\frac{1}{2})a$ (we use real distances, not the dimensionless distances for comparison with experimental data) $P(r_{k})={n_{k}}/{\sum n_{k}}$, where $n_{k}$ is the number of $r_{ij}$ values that fall within box $k$. The least-squares distance $S$ between $P(r)$ and $P_{\mathrm{exp}}(r)$, the normalised distribution function obtained from experimental data, interpolated to get its values at points $r_{k}$, is given by $S=\big{[}\sum_{k}\big{(}P(r_{k})-P_{\mathrm{exp}}(r_{k})\big{)}^{2}\big{]}^{1/2}$.
In this sum only the values $r_{k}>25\;$Å are considered to take into account the limitations of the DNA model, restricted to the backbone, which cannot describe DNA at very small distances. The experimental pair-distance probability distribution function $P_{\mathrm{exp}}(r)$ contains many short-distance contributions which are not relevant for the overall conformation of the molecule. These include the distances between atoms belonging to the same base-pair for which $r\leq d\approx 25\;$Å, where $d$ is the diameter of the molecule, including bound water. Other short distances bind atoms belonging to neighbouring or next-neighbouring base pairs, which are also within $r\lesssim 25\;$Å. The model, which represents one base-pair by a single point does not include all those contributions, and even slightly above this threshold it tends to underestimate $P_{\mathrm{exp}}(r)$. On the other hand, for larger distances, the backbone gives a meaningful picture of the conformation of a DNA molecule and its flexibility, which is the aim of our analysis. If $S$ is smaller than a predefined threshold $S_{0}$ the configuration is saved in a file for further analysis. To make sure that we only keep the conformations that provide an optimal fit to our experimental data, $S_{0}$ is selected so that only $10$ to $50$ conformations are saved for one million generated conformers. To ensure the thorough exploration of the conformational space of samples a calculation can generate up to $5\,10^{8}$ trial conformations, which is possible even with modest computing facilities because the model is sufficiently simple, with only two parameters per DNA base pair. The same analysis with an all-atom DNA model would be beyond even the best computing facilities.
2.5.3 Analysis of the saved conformations
Among the saved conformations, we select the 1000 which have the smallest least square difference $S$ between $P(r)$ and $P_{\mathrm{exp}}(r)$ and we use these “best” conformations for statistical analysis. An individual conformation has little physical meaning because at room temperature a 60 bp DNA molecule is sufficiently flexible to allow its shape to fluctuate widely. Nevertheless its properties are reflected in the statistics of these fluctuations.
In the context of protein binding, DNA curvature and flexibility are important, therefore the first relevant set of statistics concern the bending angles $\theta_{n}(i)$ where $n$ is the index of base pairs along the sequence and $i$ ($i=1,\ldots 1000$) denotes the conformations.
To demonstrate these principles we foreshadow in Figure 4 as an example of the average over the 1000 best conformations of the absolute values $|\theta_{n}(i)|$ of the local bending angles, which measure the deviations from a straight double helix, i.e.
$\overline{|\theta_{n}(i)|}=\frac{1}{1000}\sum_{i=1}^{1000}|\theta_{n}(i)|$ (2)
as well as their standard deviations
$\sigma_{\theta}=\left[\frac{1}{1000}\sum_{i=1}^{1000}\left(|\theta_{n}(i)|-\overline{|\theta_{n}(i)|}\right)^{2}\right]^{1/2}$ (3)
We also show the average of $|\theta_{n}(i)|$ restricted to bending angles which exceed a threshold $\theta_{0}=5^{\circ}$ to stress the domains which are prone to large bending.
Figure 4 shows a typical pattern from SAXS results: it is symmetrical with respect to the centre of the sequence. This was expected because the structure factor and the probability distribution function $P_{\mathrm{exp}}(r)$ result from an average of all the possible orientations of the molecules in space. Analysing $P_{\mathrm{exp}}(r)$ in terms of conformations has already reduced this three-dimensional degeneracy and brought the data along a line corresponding to the DNA chain. However, SAXS data alone do not allow us to identify a particular end of the sequence. To discuss the results we would like to compare the sequence with its base pairs identified by an index $p$, $p=1\ldots 60$, and the angles along the model $|\theta_{n}|$, $n=1,\ldots 60$, but we face a dilemma that the experiment cannot solve: it cannot tell us whether $p=1$ corresponds to $n=1$ or whether it corresponds to $n=60$. The two peaks may be real, or one of them may be an artefact because the statistical analysis mixes sequences with $n=p$ and others with n=60p+1n=60-p+1 so that we sum up data with a single peak on the right and data with a single peak on the left.
Figure 4: Statistics of the local bending angles $|\theta_{n}(i)|$ over the 1000 best saved conformations for the GAGE6 sample: the red full line shows $\overline{|\theta_{n}(i)|}$. The black line with circles shows the same average, limited to the bending angles which are above $\theta_{0}=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$
Table 1: Small Angle X-ray scattering data collection parameters
(A) Sample details GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
Organism Homo Sapiens Variant of original GAGE6 sequence Variant of original GAGE6 sequence Variant of original GAGE6 sequence
Source (Catalogue No. or reference) doi:10.1021 /acs.biochem.2c00192.
Scattering particle composition
DNA/RNA(s) 60 bp dsDNA 60 bp dsDNA 60 bp dsDNA 60 bp dsDNA
Stoichiometry of components Single component Single component Single component Single component
Sample environment /configuration
Solvent composition 150 mM Kcl, 20 mM HEPES (pH 7.4), 1 mM DTT, 5% glycerol and 5 mM MgCl2 150 mM Kcl, 20 mM HEPES (pH 7.4), 1 mM DTT, 5% glycerol and 5 mM MgCl2 150 mM Kcl, 20 mM HEPES (pH 7.4), 1 mM DTT, 5% glycerol and 5 mM MgCl2 150 mM Kcl, 20 mM HEPES (pH 7.4), 1 mM DTT, 5% glycerol and 5 mM MgCl2
Sample temperature (C) $25^{\circ}$C $25^{\circ}$C $25^{\circ}$C $25^{\circ}$C
In beam sample cell Co-flow Co-flow Co-flow Co-flow
Size Exclusion Chromatograpy SEC-SAS
[…]
Sample injection volume, mL 0.05 0.05 0.05 0.05
SEC column type S200 5/150 S200 5/150 S200 5/150 S200 5/150
SEC flow rate, mL/min 0.4 0.4 0.4 0.4
(B) SAS Data Collection GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
Data acquisition/reduction software Scatterbrain 2.82 Scatterbrain 2.82 Scatterbrain 2.82 Scatterbrain 2.82
Source/instrument description or reference Australian Synchrotron SAXS/WAXS beamline Australian Synchrotron SAXS/WAXS beamline Australian Synchrotron SAXS/WAXS beamline Australian Synchrotron SAXS/WAXS beamline
Wavelength (nm) 0.10781 0.10781 0.10781 0.10781
Camera length (mm) 3000 2385 2385 2385
Measured $q$-range ($q_{min}-q_{max})\,$Å-1 0.00450.490.0045-0.49 0.00680.620.0068-0.62 0.00680.620.0068-0.62 0.00680.620.0068-0.62
Method for scaling intensities Absolute scaling against water Absolute scaling against water Absolute scaling against water Absolute scaling against water
Exposure time(s), number of exposures. frames 183194183-194 selected (12 1 second frames) frames 156162156-162 selected (7 1 second frames) frames 163164163-164 selected (2 1 second frames) frames 166172166-172 selected (7 1 second frames)
(C) SAS-derived structural parameters GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
Methods/Software ATSAS 4.0 ATSAS 4.0 ATSAS 4.0 ATSAS 4.0
Guinier Analysis
$I(0)\pm s$ (cm-1 a.u) $005\pm 0.00013 0.0$$15\pm 0.00026 0.0$$13\pm 0.00028 0.0$$016\pm 0.00018$
$R_{g}\pm s\,$(Å) $49.89\pm 3.55$ $60.22\pm 2.23$ $54.21\pm 2.55$ $56.37\pm 1.52$
$\mathrm{min}\leq qR_{g}\leq\mathrm{max}$ limit (or data point range) 0.370.960.37-0.96 0.521.010.52-1.01 0.421.060.42-1.06 0.3810.38-1
Linear fit assessment (fidelity in PRIMUS) 1 0.71 0.9 0.97
Point range 5225-22 5225-22 5225-22 5225-22
$\mathbf{PDDF/P(r)}$ analysis ATSAS 3.2.1 ATSAS 3.2.1 ATSAS 3.2.1 ATSAS 3.2.1
$I(0)\pm s$ (cm-1 a.u) $00509\pm 0.0001516 0.0$$1459\pm 0.0001741 0.0$$1250\pm 0.0002808 0.0$$01554\pm 0.0002117$
$R_{g}\pm s\,$(Å) $55.18\pm 1.829$ $4.24\pm 1.031 55$$37\pm 1.860 55.$$5.90\pm 1.380$
$d_{\mathrm{max}}\,$(Å) 195 192 184.7 215
$q$-range (Å-1) 0.00800.17040.0080-0.1704 0.01040.15540.0104-0.1554 0.00770.15180.0077-0.1518 0.00680.15360.0068-0.1536
$P(r)$ fit assessment (total quality estimate) 0.67 (reasonable) 0.70 (reasonable) 0.67 (reasonable) 0.68 (reasonable)
Alpha 1.308 0.552 0.413 1.48
Table 1 continued: Small Angle X-ray scattering data collection parameters
(D) Scattering particle size GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
Methods/Software ATSAS 3.2.1 ATSAS 3.2.1 ATSAS 3.2.1 ATSAS 3.2.1
Volume estimates
Volume (Å-3) (MoW method) 43073 69629 69523 79329
Molecular weight (M) estimates (kDa)
From chemical composition 37.101 37.105 37.111 37.116
From SAS concentration independent method (volume of correlation method) 30.06 31.61 35.53 34.95
(D) Scattering particle size GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
(E) Data deposition
SASBDB ID: SASDT72 SASDV87 SASDV97 SASDVA7
However, if the sequence has some intrinsic asymmetry this gives an additional piece of information which can be combined with the SAXS data to solve the dilemma by examining the conformations one by one and reversing some of them so that the peak of large $|\theta_{n}|$ is always on the same side, for instance to the right that we call “forward” conformation. The algorithm to achieve what we call the “orientation” of the conformations is the following. We identify the indices $n_{1}$, $n_{2}$ on both sides of the large $n$ peak, for instance $36$ and $48$ in the case of the GAGE6 sample of Fig. 4. The mirror positions are $n^{\prime}_{1}=N-n_{2}+1$ and $n^{\prime}_{2}=N-n_{1}+1$. Then we scan again the 1000 best conformations and compute for each of them $s=\sum|\theta_{n}|,n=n_{1}\ldots n_{2}$ and $s^{\prime}=\sum|\theta_{n}|,n=n^{\prime}_{1}\ldots n^{\prime}_{2}$. If the difference between $s$ and $s^{\prime}$ is significant, then one of the two peaks is ficticious. If $s>s^{\prime}$ this conformation is already in the forward orientation. If $s^{\prime}>s$ the conformation has to be reversed by switching indices $n$ into $n^{\prime}=N-n+1$. After this check, constructing a plot similar to Fig. 4 with all conformations oriented “forward”, only one of the two peaks remains. As shown in the Results section, this is what happens for the 4 samples that we studied. When reversing a conformation, we also reverse the indices for the dihedral angles $\phi_{n}$ paying attention to the fact that, while $n$ varies from $2$ to N1N-1 for $\theta_{n}$, it varies from $2$ to N2N-2 for $\phi_{n}$ so that, instead of $n^{\prime}$ one has to use $n^{\prime\prime}=N-n$ for $\phi$. Besides the elimination of spurious signals, the orientation process improves the quality of the averaging because it becomes real averaging over the 1000 best configurations instead of the superposition of two sets of data averaged over about 500 conformations.
After this process has been completed we know that all the conformations generated by the model have the same orientation with respect to the sequence, i.e. that for all of them $n=1$ corresponds to $p=1$ or all of them correspond to $p=60$. It does not yet tell us which of the two choices is correct, and further data processing cannot answer that. However, now that we have a reliable statistical analysis we can compare $\overline{|\theta_{n}(i)|}$ with the sequence. We can expect that the domains with large bending angles are more likely to be found in regions which are richer in AT base pairs, usually more flexible than the GC-rich regions. For instance for sequence GAGE6 (Section 3.2), after orientation we detect a single peak around sites 384438-44 (Fig. 6). In the sequence the domain p=3844p=38-44 is a large continuous domain with only AT pairs. This suggests that, for this sequence, the correct choice is $n=p$, i.e. the case for which $n=1$ corresponds to $p=1$.
This “orientation” step is important to improve the quality of the statistical analysis, but it is also essential to discuss the results, when we start to examine how the local properties of the sample compare with the local sequence. It is a step that should be included in all the SAXS studies on polymer chains. In theory, if the chains don’t have any intrinsic asymmetry, they could be tagged by an additional short domain at one end, chosen for its properties to be easily identified in the SAXS data and so allow for their orientation.
The analysis can then be extended to detect other properties of the sample. In the Results section, we show how the dihedral angles vary along the sequence, and we also make histograms $H(n,\theta)$ showing, for each site, the distribution of the bending angles among the 1000 best conformations, and similarly for $H(n,\phi)$. These figures are very helpful to get a view of the likely conformations of the samples in solution and to try to understand why some sequences may be favourable for protein binding. Because long polymers, like DNA, fluctuate extensively in solution, a statistical analysis of their most likely conformations is crucial to understanding biological phenomena. It provides information that usefully complements high-resolution structures of protein-DNA complexes, which often only show a snapshot of the final conformational state after binding.
3 Results and Discussion
The CHROMIXS chromatograms indicated the presence of one single main peak eluting in a similar position for all the duplexes, with some small shoulders on either side of the peak (Supplementary Figure S1). All eluted peaks had 260:280nm UV absorbance ratios consistent with that of pure dsDNA (Supplementary Figure S1). Each dsDNA peak had some instability in predicted $R_{g}$ values across the eluted peak, possibly because of neighbouring peaks. However, frames were selected for analysis from regions where the $R_{g}$ was stable across consecutive frames within approximately 55\,Å or less (Supplementary Figure S1). In the case of GAGE6, the $R_{g}$ across the chosen frames varied from $45.3$ to 50.7850.78\,Å with an average of 48.0848.08\,Å. For GAGE6_1 the $R_{g}$ varied from 49.949.9\,Å to 52.952.9\,Å  with an average of 52.052.0\,Å. For GAGE6_2 the $R_{g}$ of the chosen frames ranged from 53.954.053.9-54.0\,Å with the $R_{g}$ in the surrounding frames varying from 51.554.051.5-54.0\,Å. In the case of GAGE6_3, the variation in the predicted $R_{g}$ across the chosen frames was minimal, ranging from 49.350.649.3-50.6\,Å  with an average of 50.150.1\,Å.
All scattering shown as $\log_{10}(I)$ vs. $\log_{10}(q)$ plots demonstrated good quality data with a plateau in the Guinier region (Fig. 5). When constrained to $q.Rg\lesssim 1$ (suitable for elongated molecules), the Guinier fits had high fidelity scores and acceptable distributions of residuals. The fits also provided $R_{g}$ values similar to the expected value for a straight 60bp dsDNA duplex as simulated with CRYSOL ($R_{g}=56.10\,$Å) (Fig. 5). The distance distribution functions of the various oligonucleotides from this study highlight some robust features across each variant. Depending on the duplex the shape of the $P_{\mathrm{exp}}(r)$ function changes, with characteristic bumps in the GAGE6, GAGE6_1, and GAGE6_2 functions which then almost disappear in GAGE6_3 as shown in Figs. 6, 7, 8, 9 and 10 in the next sections. The $D_{\mathrm{max}}$ values for all functions were within the range of ±10% of the theoretical $D_{\mathrm{max}}$ of a 60 bp duplex. Supplementary Figure S4 shows that deviation of $D_{\mathrm{max}}$ from the optimal value given by our procedure only leads to small variations in $P_{\mathrm{exp}}(r)$. Additionally, Supplementary Figure S5 shows that despite variation of the GNOM regularisation parameter the bumps in our respective $P_{\mathrm{exp}}(r)$ functions persist. The $R_{g}$ values derived from the individual $P_{\mathrm{exp}}(r)$ functions were more or less consistent with those from the Guinier approximation. We noted minor inconsistencies between $P_{\mathrm{exp}}(r)$ and Guinier-derived $R_{g}$ for GAGE6 and GAGE6_1. The details concerning SAXS parameters are summarised in Table 1.
3.1 Small Angle X-ray Scattering data
Figure 5: Scattering data and Guinier fits of the various GAGE6 oligonucleotides: $I(q)$ versus $q$ plot, Guinier plot ($\ln[I(q)]$ vs $q^{2}$), and Guinier residual plots for (A) GAGE6, (B) GAGE6_1, (C) GAGE6_2, (D) GAGE6_3. Guinier fit parameters and results are tabulated on the figure.
3.2 Sequence GAGE6
Let us start with the GAGE6 sequence, which is the leading sequence of the series since it is a native promoter sequence that is bound by SFPQ. Figure 6A compares the experimental pair-distance distribution function $P_{\mathrm{exp}}(r)$ with $P(r)$ calculated from the 1000 best conformations saved after running Monte-Carlo simulations to generate conformations with the polymer model. The agreement between the two distributions is evaluated by computing
$\chi^{2}=\frac{1}{N_{r}-1}\sum_{1}^{N_{r}}\left[\frac{P(r_{j})-P_{\mathrm{exp}}(r_{j})}{\sigma_{j}}\right]^{2}\;,$ (4)
where $r_{j}$ are the points where $P_{\mathrm{exp}}(r)$ has been calculated, $\sigma_{j}$ the experimental error at these points (estimated by the GNOM program), and $N_{r}$ the number of calculation points. The summation is restricted to $r_{j}>25\;$Å because the polymer model cannot describe the internal structure of DNA, which dominates $P_{\mathrm{exp}}(r)$ for r<25r<25\;Å as discussed above (section 2.5.2). The curve deduced from the selected conformations of the polymer model (Figure 6A) provides a very good description of the experimental curve in the whole domain where the polymer model is valid, indicating in this instance that the polymer model ensemble can very accurately reproduce $P_{\mathrm{exp}}(r)$.
Figure 6: Sequence GAGE6: Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve). The grey region, for r<25r<25\,Å  is the low-$r$ domain in which the model cannot describe the internal structure of the double helix. The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6: Statistics of the local bending angles $\theta_{n}(i)$ and dihedral angles $\phi_{n}(i)$ over the 1000 best-saved conformations for the GAGE6 sample, after orientation of the conformations as discussed in the Materials and Methods section. The red full line shows the average over $i$ of $|\theta_{n}(i)|$. The black line with circles shows the same average, limited to the bending angles which are above $\theta=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. The thin black line, without symbols, shows the average over $i$ of the dihedral angles and the thin black dash-dot line shows their standard deviation.(C) Sequence GAGE6: Three-dimensional plot of the normalised probability distribution function $H(n,\theta)$. The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$. The value $W$ marked on the figure shows the fraction of conformations which have a bending angle above the value shown below $W$ within the domain indicated by the two values of $n$.
Figure 6B shows the statistics of the bending angles $|\theta_{n}|$ and dihedral angles $\phi_{n}$ with their averages over conformations and standard deviations. For the bending angles this figure is analogous to Fig. 4 except that the conformations have been oriented as described in the Materials and Methods section. Only one of the double maxima that were forming a mirror image with respect to the middle of the sequence has subsisted, which confirms that the symmetric pattern was created through symmetry inherent in the SAXS measurement. The value $\Delta s/s$ marked on the figure corresponds to $(s^{\prime}-s)/s$, where $s$ and $s^{\prime}$ were the local sums of the bending angles in the two mirror domains used to orient the conformations. Their differences were significant, suggesting that one of the two mirror images was fictitious, which is confirmed by the figure as one of them has been fully erased in the orientation step.
The three-dimensional plot of Fig. 6C shows the probability distribution of the bending angles $|\theta|$ for each site $n$ along the sequence. These distributions are normalised
$\int_{0}^{\pi}H(n,\theta)=1\quad{\mathrm{for~all~}}n$ (5)
which allows us to calculate the fraction $W$ of conformations which have a bending angle above some value in a given domain. The result $W=0.06557$ is shown in Fig 6C for $35\leq n\leq 49$ and a lower value of $\theta$ of $45^{\circ}$. Thus, although most of the bending angles are well below $30^{\circ}$, in some regions of the sequence there are about $6\%$ of the conformations which show bending above $45^{\circ}$. The statistics clearly show that one region of the sequence is special, with a much higher flexibility and larger bending than elsewhere. The colour view of the three-dimensional histogram Fig. 6C, (see Supplementary material Fig. S8) suggests that the large bending angles for $35\leq n\leq 49$ are separated from the bulk of smaller bending angles, as if they result from some permanent bend in this region and not simply from large amplitude thermal fluctuations. As this part of the histogram only concerns $6\%$ of the conformations, the resolution of the analysis of the SAXS data does not however allow us to make a definitive statement on this point.
We would like to stress that these statistical results which provide precise information on the local properties of the sample, are extracted from a probability distribution function which looks rather featureless although it results from a large ensemble of pair distances. It is remarkable that this smooth looking, almost straight function, actually encodes so much information. Now that we have oriented the conformations we can also compare the statistical results with the actual sequence of the sample. There is still an unknown point: should the sequence listed in Fig. 1 be oriented from 1 to 60 or from 60 to 1 ?
The GAGE6 sequence is asymmetric, with large AT-rich domains on the 3’ end, while the 5’ end is more GC-rich. The AT base pairs, linked by only two hydrogen bonds, are generally more flexible than the GC pairs linked by three hydrogen bonds because their weaker bonding allows greater freedom in their local conformations. This effect is also enhanced by easier fluctuational openings, even at room temperature [DNAFLEX]. Therefore it seems reasonable to assume that the $n=60$ side of our diagrams corresponds to the 3’ end. With this assignment, we can notice that the domain $n=35\ldots 47$ that we used for the orientation step is almost completely made of AT pairs with only two, separate, GC pairs inside it. Therefore it is not surprising to find that this domain is highly flexible as found in the analysis of the SAXS data. Thus the results of our analysis of the SAXS data suggest that this polymer modelling pipeline is able to detect specific features at the scale of a few base pairs. Its validity is strengthened by the results on the fluctuations of the dihedral angles, which show a maximum in their standard deviations which coincides with the maximum of the bending angles, although it is slightly broader. In the linear chain of joined segments that we use to analyse the data there is no geometrical constraint which links bending and twist. But in DNA there is such a constraint due to the double helical structure. It is easy to check with a mechanical device such as two wires twisted together to make a helix that a large amount of bending tends to locally untwist the device, simply by geometrical effects. The fit of our polymer model to $P_{\mathrm{exp}}(r)$ detects this phenomenon although it is not built into our hypothesis.
However, although the sequence strongly contributes to determining the local flexibility of the sample the results also point to subtle effects associated with some cooperativity along the double helix. In the GAGE6 sequence, sites 384438-44 make a large continuous domain with only AT pairs and therefore it is not surprising to find higher bending angles in this region as pointed out above. But the GAGE6 sequence also has a 4-AT-pairs domain at sites 283128-31 which are not within the region where we detect larger bending angles. Similarly, the 5-AT-pairs domain at sites 505450-54 also show fairly low bending angles.
The sequence gives a hint on the local flexibility of DNA but the analysis of the SAXS data suggests that experiments which are sensitive to the conformation of DNA tell us more. This may have many biological implications. Within cells and viruses, DNA has to be tightly bent or kinked at many sites to allow for compact packaging of large amounts of genetic material [WIDOM1984]. Many studies have tried to address the question of whether DNA bending arises from metastable bent structures or more flexible “joints” in dsDNA. Previously solved structures of dsDNA AT-rich dodecamers [STEFL2004, HIZVER2001] suggest that bent duplexes are stable and are maintained in a bent conformation by several molecular features. This notion was supported by a fluorescence polarisation experiment [CHIRICO2001] that also argued for a permanent static bend in some AT-rich dsDNA sequences. A review by Harteis and Schneider [HARTEIS2014] on the structure of DNA summarised that whilst AT tracts can introduce an intrinsic bend, they in fact somewhat increase the rigidity of such sequences due to certain stabilising effects that accommodate the bending of DNA: propellor twisting of bases, narrow minor groove, wider major groove [STEFL2004, HIZVER2001] as well as non-Watson-Crick hydrogen bonds forming down the major groove. Rather interestingly, the bending of AT-rich DNA was also shown to be sensitive to the order of the bases, with A4T4 vs. T4A4 steps having different effects, the former causing a global bend in the helix, and the latter forming a straighter helix [STEFL2004]. DNA has been considered to act more flexibly when a “kink” defined as the local unstacking of base pairs causing a change in the orientation of the helix appears [HARTEIS2014]. Pyrmidine-purine (TA; CA) steps reportedly exhibit the lowest base-stacking energy and can behave more like flexible hinges that can cause DNA bending.
In addition to these structural effects, further studies have also examined the fluctuational opening of DNA base pairs or the formation of “bubbles” and whether these could be a source of flexibility in DNA. Gueron et al. [GUERON1987] described the probability of the fluctuational opening of base pairs as low via imino-proton exchange. More recently Theodorakoupolous and Peyrard [DNAFLEX] reviewed this concept and theorised that below the melting point of a duplex, an increase in temperature towards physiological temperature may increase the opening of base pairs enough to act as flexible hinges in dsDNA, which may have wider implications for DNA structure and recognition by proteins.
Figure 7: Sequence GAGE6_1: Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve). The grey region, for r<25r<25\,Å  is the low-$r$ domain in which the model cannot describe the internal structure of the double helix. The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_1: Statistics of the local bending angles $\theta_{n}(i)$ and dihedral angles $\phi_{n}(i)$ over the 1000 best-saved conformations for the GAGE6_1 sample, after orientation of the conformations as discussed in the Materials and Methods section. The red full line shows the average over $i$ of $|\theta_{n}(i)|$. The black line with circles shows the same average, limited to the bending angles which are above $\theta=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. The thin black line, without symbols, shows the average over $i$ of the dihedral angles and the thin black dash-dot line shows their standard deviation.(C) Sequence GAGE6_1: Three-dimensional plot of the normalised probability distribution function $H(n,\theta)$. The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$. The value $W$ marked on the figure shows the fraction of conformations which have a bending angle above the value shown below $W$ within the domain indicated by the two values of $n$.
Alongside the fluctuational opening of the duplex, further unwinding of DNA regions is also important. Taking a simple physical model of 2 interwound strands and bending it sufficiently causes unwinding of the 2 strands at the point of bending. Recently, Chandrashekhar et al. [CHANDRASEKAR2024] demonstrated via a minicircle assay that DNA can locally unwind in response to bending. This has implications for gene regulatory and DNA packaging elements, as the bending of DNA by any of the aforementioned mechanisms may also contribute to the unwinding of the duplex.
In a SAXS experiment, it can be complicated to delineate between interrelated effects such as stable, rigid/bent polyAT structures, flexibly kinked DNA as a result of unstacked base pairs, the unwinding of strands, or the fluctuational opening of base pairs. This is particularly true given that SAXS is relatively speaking a low-resolution structural technique. However, having shown that we can quantitatively detect the bending of DNA in solution we hypothesised that such events could be described by the complementary study of a series of related sequences derived from GAGE6.
3.3 Sequences GAGE6_1, GAGE6_2, GAGE6_3
This series of three sequences was derived from GAGE6 by incrementally removing its longest AT-rich domains to replace them with GC-rich segments. This allowed us to test two things
  • (i)
    to check how the analysis of the SAXS data detects these local changes and how they influence the conformations of the samples, and
  • (ii)
    to try and understand the properties of the GAGE6 sequence that allow it to bind SFPQ. The first sequence of the series, GAGE6_1, only differs from GAGE6 by four bases in sites 283128-31. The TTTT sequence in GAGE6 has been replaced by GCGC in GAGE6_1.
Figure 7 shows the equivalent data for the GAGE6_1 variant to Figure 6 for GAGE6. Figure 7A shows that, although the change in the sequence is very localised, the shape of $P_{\mathrm{exp}}(r)$ is significantly modified: the single bump around 100100\,Å in GAGE6 splits into two lumps around $70$ and 125125\,Å  for GAGE6_1. Our analysis shows that this difference is due to a large change in the distribution of $\theta$ angles (Fig. 7B). The broad domain with large $\theta$ angles covering sites 364636-46 in GAGE6 is now narrower, and restricted to sites 445244-52 only, but is also much sharper as the maximum of $\theta_{n}(i)$ reaches $36^{\circ}$ instead of about $20^{\circ}$ for GAGE6. The dihedral angles show the same behaviour, with large twist fluctuations being restricted to the same narrow domain.
Figure 7C confirms the presence of large bending angles in a narrow domain and the integration of $H(n,\theta)$ for $41\leq n\leq 52$ and $\theta>45^{\circ}$ shows that the fraction $W$ of bending angles above $45^{\circ}$ has drastically increased from 6.4% to 12.6%. The GAGE6_1 sequence therefore has a high probability of being sharply bent in a narrow region. It is interesting to notice the qualitative analogy between the experimental pair-distance distribution function for GAGE6_1 and the double-bump simulated $P(r)$ (Fig. 2) obtained with an atomistic model of sharply bent DNA. This strengthens the notion that $P(r)$ has a high sensitivity to molecular conformations, particularly bending, with sharp bending generating a function with two bumps.
Figure 8: Sequence GAGE6_2: Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve). The grey region, for r<25r<25\,Å  is the low-$r$ domain in which the model cannot describe the internal structure of the double helix. The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_2: Statistics of the local bending angles $\theta_{n}(i)$ and dihedral angles $\phi_{n}(i)$ over the 1000 best-saved conformations for the GAGE6_2 sample, after orientation of the conformations as discussed in the Materials and Methods section. The red full line shows the average over $i$ of $|\theta_{n}(i)|$. The black line with circles shows the same average, limited to the bending angles which are above $\theta=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. The thin black line, without symbols, shows the average over $i$ of the dihedral angles and the thin black dash-dot line shows their standard deviation.(C) Sequence GAGE6_2: Three-dimensional plot of the normalised probability distribution function $H(n,\theta)$. The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$. The value $W$ marked on the figure shows the fraction of conformations which have a bending angle above the value shown below $W$ within the domain indicated by the two values of $n$.
In GAGE6_2, in addition to replacing the AT base pairs between positions $28\leq n\leq 31$ with GC pairs, the AT pairs in positions $38\leq n\leq 44$ have now also been replaced with GC pairs. As expected, the change in properties of the sample with respect to GAGE6 is even more drastic. Figure 8A shows that $P_{\mathrm{exp}}(r)$ has sharper lumps, slightly moved towards larger values of $r$. The error bars given by GNOM in the calculation of $P_{\mathrm{exp}}(r)$ from $I(q)$ are larger and the fit of this pair-distance distribution function by the simple model is less accurate although it remains within the error bars almost everywhere.
Figure 8B shows that the distribution of large bending angles along the sequence, as well as the domains where the dihedral angles have large standard deviations, have qualitatively changed. Figure 8B appears to be very consistent with the sequence of GAGE6_2. Its largest AT domains are $10\leq n\leq 13$ and a large domain from $n=50$ to $n=57$ which is only interrupted by a single GC pair at position $n=55$. This agrees well with the small increase of bending and twist fluctuations seen around $n=10$ in Fig. 8B and the larger bending angles and twist fluctuations in the range $44\leq n\leq 57$ split into two peaks by a dip in sites $52,53$ which could correspond to the GC pair at site $55$ within the resolution given by the analysis of the SAXS data. The bending angles $|\theta_{n}(i)|$ in this domain are however much smaller than in the sharp bending domain of GAGE6_1. This qualitative change from GAGE6_1 may nevertheless seem surprising if we think that the only difference in sequence between GAGE6_1 and GAGE6_2 is the elimination of an AT-rich domain between positions $38\leq n\leq 44$. Both GAGE6_1 and GAGE6_2 have a large AT-rich domain ($50\leq n\leq 57$) but in GAGE6_2 this region causes a smaller bending effect than GAGE6_1. This points out again that, although sequence and flexibility are correlated, the preferred conformations of a DNA sequence in solution are not only determined by the local sequence but collective effects at larger distances are also important. The histograms of the bending angles shown in Fig. 8C and Supplementary Fig. S8 are somewhat similar to the data from Fig. 7: the fraction of large bending angles in the range $44\leq n\leq 57$ is of the order of 9.6%, suggesting that GAGE6_2 is also sharply bent in a narrow region.
Figure 9: Top panel: sequence GAGE6_2. Comparison of AutoGNOM-4-$P_{\mathrm{exp}}(r)$ (small circle points and thin black curve with error bars) and $P(r)$ given by the 1000 best conformations of the polymer model (thick red curve). The gray region for r<25r<25\;Å is the low-r domain in which the model cannot describe the internal structure of the double helix. Bottom panel: Statistics of the local bending angles $\theta_{n}(i)$ over the 1000 best model conformations: the red full line shows the average over $i$ of $|\theta_{n}(i)|$. The thin black line without symbols shows the same average limited to the bending angles $\theta$ above $5^{\circ}$ and dashed blue curve shows the standard deviations of $\theta_{n}(i)$. The thin black line without symbols shows the average of the dihedral angles and the thin black-dot-line shows their standard deviation. The vertical dashed lines show the boundaries of the domain where large bending was observed in the analysis of AutoGNOM-5-$P_{\mathrm{exp}}(r)$.
However, the results shown in Fig. 8A raise questions because the fit of the polymer model to the $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 looks rather poor. We know that the model is oversimplified to fully describe DNA, but, for r>25r>25\;Å it should nevertheless provide a better description of the experiment because we have extensively explored the conformational space of the sample by generating more than $4\,10^{8}$ conformations and selecting the best 1000 for the analysis. This failure suggests that the difficulty could come from $P_{\mathrm{exp}}(r)$ itself. At this point we would like to stress the importance of choosing a model which has realistic bending and torsional properties to make sure that the conformations which enter into our fits are actually accessible to DNA molecules. A purely random model, such as the Gaussian model which can adopt any conformation at no energy cost might be able to provide a good fit to $P_{\mathrm{exp}}(r)$, but with unphysical DNA conformations. In section 2.4 we discussed that, for sequence GAGE6_2, the determination of the optimal regularisation parameter $\alpha$ by GNOM is particularly hard. GNOM-5 converges to $\alpha=0.01084$ while GNOM-4 converges to $\alpha=1.31$ and generates a much smoother $P_{\mathrm{exp}}(r)$ with smaller estimated error bars. We conjectured that, for a given sample, the functions $P_{\mathrm{exp}}(r)$ for different values of $\alpha$ could contain almost the same information on the physical properties of the DNA sample. To check this conjecture we repeated with the AutoGNOM-4-$P_{\mathrm{exp}}(r)$ functions the same analysis initially performed with those derived using AutoGNOM-5. The results are shown on Fig. 9.
The top panel of Fig. 9 shows that the model gives an almost perfect fit of AutoGNOM-4-$P_{\mathrm{exp}}(r)$ which is much smoother than AutoGNOM-5-$P_{\mathrm{exp}}(r)$. In comparing Fig. 8 and Fig. 9, one should keep in mind that the definition of $\chi^{2}$ (Eq. (6)) compares the discrepancy between $P_{\mathrm{exp}}(r)$ and the theoretical $P(r)$ with the estimated error on $P_{\mathrm{exp}}(r)$. As the error bars on AutoGNOM-4-$P_{\mathrm{exp}}(r)$ are much smaller than on AutoGNOM-5-$P_{\mathrm{exp}}(r)$ similar values for $\chi^{2}$ in both figures imply that the discrepancies are much smaller in Fig. 9. The bottom panel of Fig. 9 shows that, although the two functions $P_{\mathrm{exp}}(r)$ look radically different, they provide almost the same results concerning the propensity of the DNA sample to bend in the 44-57 domain which is the AT-rich domain. However, the AutoGNOM-4-$P_{\mathrm{exp}}(r)$ misses the large degree of bending occurring near n=5657n=56-57 which is detected by the AutoGNOM-5-$P_{\mathrm{exp}}(r)$. Although it cannot be formally excluded that the bending detected by AutoGNOM-5-$P_{\mathrm{exp}}(r)$ in the vicinity of an AT pair is an artefact, it is likely that the high-value of $\alpha$ derived by AutoGNOM-4 leads to the over-regularisation of this function, which smoothes out this local effect. Nevertheless it is remarkable that two functions $P_{\mathrm{exp}}(r)$ which look so different at a first glance encode the same information about the bendability of a sample near one of its ends. This supports our conjecture and implies that a perfect refinement of $P_{\mathrm{exp}}(r)$, which may be difficult, is not an absolute requirement to analyse SAXS data with a polymer model as proposed in this work.
In GAGE6_3, the last of the series of sequences derived from GAGE6, the last large AT-rich domain that spanned positions $50\leq n\leq 54$ has been replaced entirely by GC base pairs. Figure 10A shows that $P_{\mathrm{exp}}(r)$ has lost its two-lump structure and conforms in comparison to the other experimental functions more to the simulated shape of a straight duplex (Figure 2). The disagreement between the polymer model fit and $P_{\mathrm{exp}}(r)$ at large values of $r$ here may be a result of the former not accounting for hydration shell effects whilst the latter does. Although the large AT-rich regions of the $n>30$ range that we used to orient the conformations for previous sequences are no longer present, the sequence still has asymmetry. The domain $10\leq n\leq 26$ has several short AT sequences while its mirror image $35\leq n\leq 49$ now only has three AT pairs and is likely to be less flexible, being systematically enriched for GC content. This is confirmed by the sum $s$ and $s^{\prime}$ of $|\theta_{n}|$ in these two regions which differ systematically from each other and can be used to orient the conformations with respect to the sequence. Figure 10B, which does not display the symmetry expected for unoriented SAXS data, shows that the orientation procedure described in the Material and Methods section is again valid for GAGE6_3. Figure 10B shows a broad domain with large bending angles with a sharp maximum at position $n=26$ which coincides with two AT pairs in the middle of a GC sequence. The three-dimensional plot of the histograms $H(n,\theta)$ plotted in Fig. 10C and Supplementary Fig. S8 shows larger bending angles in the domain $17\leq n\leq 35$ but the integrated weight of the histograms for angles exceeding $45^{\circ}$ in this region shows that only 3.8% of the conformations exhibit large bending. As suggested by its sequence, mostly GC base pairs and only a few small domains with AT pairs, not wider than three consecutive sites, the GAGE6_3 sequence has mostly straight conformations.
Figure 10: Sequence GAGE6_3: Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve). The grey region, for r<25r<25\,Å  is the low-$r$ domain in which the model cannot describe the internal structure of the double helix. The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_3: Statistics of the local bending angles $\theta_{n}(i)$ and dihedral angles $\phi_{n}(i)$ over the 1000 best-saved conformations for the GAGE6_3 sample, after orientation of the conformations as discussed in the Materials and Methods section. The red full line shows the average over $i$ of $|\theta_{n}(i)|$. The black line with circles shows the same average, limited to the bending angles which are above $\theta=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. The thin black line, without symbols, shows the average over $i$ of the dihedral angles and the thin black dash-dot line shows their standard deviation.(C) Sequence GAGE6_3: Three-dimensional plot of the normalised probability distribution function $H(n,\theta)$. The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$. The value $W$ marked on the figure shows the fraction of conformations which have a bending angle above the value shown below $W$ within the domain indicated by the two values of $n$.
Supplementary figure S9 shows the results of the analysis of the functions $P_{\mathrm{exp}}(r)$ obtained with Auto-GNOM-4 for the four samples (see supplementary figure Fig. S5). It can be compared with Figs. 6, 7, 9 and 10 showing the results obtained with Auto-GNOM-5. It confirms our conclusions about the domains prone to extra flexibility and bending in each sequence. However, it appears that large regularisation parameters $\alpha$, giving smoother $P_{\mathrm{exp}}(r)$, may lead to lower amplitudes of the calculated bending and flexibility in the DNA sites belonging to those domains, and, in some cases loss of fine structures. Our results suggest that, for an analysis with a model which properly describes DNA bending and torsional rigidity, and therefore avoids unphysical conformations, as the choice of the optimal regularisation parameter is difficult, favouring small regularisation parameters $\alpha$ may be a good choice because it leads to $P_{\mathrm{exp}}(r)$ functions which encode more structural details.
The study of a series of four DNA sequences, including GAGE6 and three others obtained by gradually replacing AT-rich domains with GC tracts is interesting because the unique structural features of bent AT-tract-containing duplexes can be important for the recognition of dsDNA by proteins [HARTEIS2014]. Typically, the major groove in dsDNA has the highest potential for protein-base recognition as the functional groups of the four bases are accessible [HARTEIS2014]. In the case of bent AT-tract-containing structures, the major groove shows an even higher potential for base recognition as the groove is often widened by bending [HARTEIS2014, STEFL2004]. Bending likely serves to enhance sequence-specific recognition of DNA by proteins in many contexts. A simultaneous effect in AT-tract containing dsDNA is the narrowing of the minor groove [STEFL2004]. The narrower groove places elements of the backbone much closer to one another resulting in an enhanced electrostatic effect that can facilitate structure-specific rather than sequence-specific recognition by proteins [FERRARI1992, ROHS2009].
It is possible that SFPQ recognises the GAGE6 oligonucleotide through an interaction of the RGG domains with either a wider major groove or a more pronounced electrostatic interaction with the backbone due to the narrower minor groove. The latter is an interesting possibility given that RGG motifs are the putative DNA-binding regions of SFPQ, and arginine is the most common residue found to interact with narrow minor grooves [ROHS2009]. Another possibility is that the fluctuational opening of the “TTTATTT” region in the GAGE6 oligo, which was the approximate region where the majority of bending was detected in our analysis, plays a role in protein binding. This effect may be more pronounced at a higher temperature [DNAFLEX] or inside biological condensates which through multivalent competitive interactions can reportedly cause the melting of dsDNA [NOTT2016]. The truncation of the GAGE6 sequence from 60 bp to 40 bp reduced the binding of SFPQ from 84% to 45% as calculated via a gel shift assay [WANG2022]. This truncation resulted in the “TTTATTT” domain, being cut in half. This suggests that the last 20bps of the sequence may provide the region required for the enhanced binding of SFPQ. However, as functional aggregation of the protein is a key component of binding [LEE, KONING2025] it is possible that 40 bp was simply too short a length to allow multiple units of SFPQ to have several points of contact with the DNA.
The idea that certain sites in dsDNA offer less resistance to structural deformation and so a lower energetic penalty to protein-DNA complexes that deform rigid DNA is also important for many protein-DNA interactions. Naturally, if a protein is required to bend DNA as part of its function, a pre-bent or flexible substrate would dramatically reduce the amount of energy required to bend DNA. This concept has been reviewed by [HARTEIS2014] and some examples are:
ECOR1 significantly bends DNA at its recognition site ‘GAATTC’ as part of cleavage reactions [WIDOM1984]. The degree of natural bending at this target site likely lowers the energy required for binding as bending of the unbound DNA can place it close to a position that already resembles the bound state. Gartenburg and Crothers [GARTENBERG1988] examined the degree of DNA bending caused by the Catabolite activator protein (CAP) in response to different mutated DNA targets. DNA bending in regions flanking the binding site enhanced the affinity of the CAP-DNA complex by creating a larger and more complex interface. HMG box domains in transcriptional regulatory proteins bind DNA. In general, all of the binding sites for these proteins are AT-rich, and the protein SRY has been shown to induce a sharp bend when binding to the sequence “AACAAG” [FERRARI1992]. It was hypothesised by Ferrari et al. [FERRARI1992] that this binding site provided a lower resistance to structural deformation and so lowered the high energetic cost of bending DNA with an already bent substrate. Nagaich et al. [NAGAICH1997] examined the role of P53 in DNA bending and deduced that P53 caused bending in a number of DNA response elements. EMSA assays indicated that increasingly bent DNA had a higher affinity interaction with P53. Whilst the examined response elements did not contain strict AT tracts, they did contain dinucleotide AT regions on either side of the response element. TATA-box binding protein (TBP) binding to its DNA target is largely mediated by the bent shape that its DNA target takes and this contributes to the affinity of the complex [HARTEIS2014]. The energetic cost of the DNA’s shape is lowered by the reduced stacking of the TA steps in the TATA sequence.
Realising the pre-bent nature of many of these DNA targets is important in understanding the affinity and energetics of certain protein-DNA interactions. Protein recognition of AT-rich DNA can be complicated because of many interrelated effects. However, DNA bending is reportedly sensitive to environmental conditions such as temperature, salt, and divalent cations [HARAN2009]. Given an initial hypothesis of unwinding, bending or fluctuational opening of dsDNA playing a role in the affinity of a protein-DNA complex it may be possible to fine-tune environmental conditions to maximise such effects and so increase the affinity of protein-DNA interactions for structural studies. Furthermore, experimental data on DNA bending in response to sequence composition may also be used to train structure prediction tools such as successors to Alphafold 3. Our polymer model approach may also perhaps present an opportunity for the development of other multiphase modelling systems, that can separately model both the DNA by itself and in complex with a protein. This could be technically possible via a contrast matching small angle neutron scattering experiment where an experimentalist could make the protein “invisible” and only analyse the DNA in an unbound and also protein-bound state.
4 Conclusion
In this work, we have shown that SAXS data can be analysed to determine the statistical properties of the conformation of short DNA sequences in solution. Although the SAXS structure factor averages over all spatial orientations of the molecules, for sequences which have some asymmetry, even when it is rather weak such as for GAGE6_3, the computed conformations can be oriented with respect to the sequence so that specific features detected by the analysis can be related to the DNA sequence.
Our analysis uses a polymer model which is simple enough to allow a very broad exploration of conformational space (up to $10^{8}$ or $10^{9}$ conformations with moderate computing facilities) but is nevertheless able to quantitatively describe the average persistence length and torsional rigidity of the DNA double helix. This is important to ensure that only conformations accessible to a DNA molecule are generated. In contrast to the protein chains in intrinsically disordered protein regions, which can be described by a random Gaussian chain [MARTIN2021], the bending and torsional rigidities of the double helix introduce strong constraints on the conformational space, which are intrinsic to our polymer model. The model, which is restricted to a chain of rigid segments, does not impose any constraints tying bending to twist. Nevertheless, our analysis detects the increase of twist fluctuations which is expected for DNA undergoing strong bending due to the mechanical properties of the double helix. The ability of the analysis to detect a property of DNA which in itself is not built into the model strengthens the credibility of the method, as well as its capability to detect regions which are prone to bending in correlation with the sequence.
As expected the results confirm that AT-rich regions are more flexible than GC-rich domains, however, our results have also shown that the link between sequence and the preferred conformations of free DNA in solution are not trivially imposed by the local sequence. Using a series of sequences derived from one another by small local changes we have exhibited unexpected properties which suggest that collective effects are also very important.
Our method relies on a fit of $P_{\mathrm{exp}}(r)$, i.e. analysis carried out in real space, rather than handling the data in reciprocal space. This introduces some difficulty because an accurate derivation of $P_{\mathrm{exp}}(r)$ from the experimental data is not straightforward although it has been the object of numerous investigations. The determination of the regularisation parameter $\alpha$ which enters into the calculation is delicate. However, our analysis has shown that, when $\alpha$ is modified, the results on the properties of the sample are robust.
It turns out that the view in real space is essential to detect local effects which are spread out in Fourier space. This is not specific to SAXS and spans various domains of physics. A famous example has been provided by nonlinear localised waves in water or solid state physics [DAUXOIS2005, SOLITONBOOK]. A numerical experiment on vibrational waves to study thermalisation in solids, by Enrico Fermi and coworkers, raised a puzzling question that was unexplained for more than 10 years. It stayed a mystery because, being a wave problem, the results were displayed in the wavevector space ($q$ space). It was only when some physicists looked in real space that they got a clue because they detected localised phenomena (later called solitons) which could easily solve the puzzle. The real and Fourier spaces bring complementary views, and although experiments and theory often lead naturally to the Fourier space, the real space approach may sometimes be the most appropriate.
This work highlights the interest of SAXS studies to investigate protein binding sites on DNA because they are able to probe the conformations of DNA in solution and their dynamics, which is reflected in the statistics of the conformations provided by SAXS. Crystallographic and CryoEM studies, widely used to determine the structure of protein-DNA complexes are very precise with their final structure, in which the conformational freedom of DNA is often highly constrained by nature of being bound to a protein or vitrification and crystallisation. However, when proteins scan DNA before binding, they often sample and test these free-DNA conformations which can be observed by SAXS studies in solution. Therefore SAXS, crystallographic, and cryoEM studies are complementary. Having shown that the analysis of SAXS data is able to determine the properties of DNA in solution at a scale of a few base pairs enhances the interest in the combination of the two methods to investigate DNA-protein complexes.
[…]
[saxsdna]
Supplementary Materials:
Figure 1: Supplementary Figure S1: Figure S1 shows the absorbance traces of all the duplexes at 214nm, 260nm and 280nm vs time (seconds) followed by scattering traces of all the duplexes plotted from CRHOMIXS showing average X-ray intensity against frame number. Predicted Rg values are shown as blue curves and were predicted using CHROMIXS. The regions used for buffer subtraction are shown in orange and the regions used for sample processing in green.
$\Delta q=0.001\,$Å-1 GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
GAGE6 - - - -
GAGE6_1 0.614 - - -
GAGE6_2 0.614 0.855 - -
GAGE6_3 0.209 0.855 0.614 -
Figure 2: Supplementary Figure S2: Comparative distance distribution functions of the different oligonucleotides used in this study. Sequence of each oligonucleotide indicated above $P_{\mathrm{exp}}(r)$ function with AT tracts highlighted in red and GC tracts in green. Sequence-dependent changes in the $P_{\mathrm{exp}}(r)$ function indicate the position of multiple bumps in the AT tract containing duplexes and a smooth descent for GAGE6_3.
Figure 3: Supplementary Figure S3: Comparative $P(r)$ functions for datasets accompanied by 1x (pink line) and 2x (black line) standard error show consistency in the shape of P(r). Errors bars only plotted for data from 1x standard error (pink).
Figure 4: Supplementary Figure S4: Variation of $D_{\mathrm{max}}$ parameter around the reported values shows that $P(r)$ features remain consistently present.
Figure 5: Supplementary Figure S5: Comparison between the functions $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (red curve with error bars) and with AutoGNOM-4 (black curve with error bars) for the four DNA samples. The values of the regularisation parameter $\alpha$ used for each curve are indicated inside each panel. The inset panels show the variation of the Total Quality Estimate (TQE) versus $\alpha$ (in logarithmic scale) calculated by AutoGNOM-4.
Figure 6: Supplementary Figure S6: Variation of the GNOM alpha parameter around the fitted value indicates that while the magnitude of features changes, their position relative to r remains consistent.
Figure 7: Supplementary Figure S7: Plots of the GNOM fits to the scattering data and their corresponding residual plots.
Figure 8: Supplementary Figure S8: Another view of the three-dimensional histograms $H(n,\theta)$ which highlights some features that are difficult to see on the three-dimensional plots. The left part shows the projection of the histograms for each site on a single plane which displays the probability of a given value of $\theta$, whatever the position in the sequence: $H_{\mathrm{sum}}(\theta)=\sum_{n=2}^{N-1}H(n,\theta)$ The right part uses a color code to show the three-dimensional histogram in a two-dimensional image. This is easier to get an idea of the map of the bending angles along the sequence.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: Supplementary figure S9: Analysis of the $P_{\mathrm{exp}}(r)$ obtained with Auto-GNOM-4 (see Fig. S5). For each sample the top panel shows the function $P_{\mathrm{exp}}(r)$ with error bars (black) and its fit by the polymer model (red full line). The bottom panel shows the statistics of the local bending and dihedral angles, as on Figs. 6, 7, 8, 10. The vertical dashed lines, limiting the domain showing the largest bending are at the same positions as on the corresponding figures of the paper. The particular case of GAGE6_2 is also shown in the paper as Fig. 9.

### v2

@@
 Figure 2: Illustration of the effect of DNA bending on the shape of $P(r)$.
-(A) A group of simple 60 bp DNA models created by Alphafold 3 [ABRAMSOL] three of which were obtained by bending the duplex at base-pair 34.
+(A) A group of simple 60 bp DNA models, the first of which was created by Alphafold 3 [ABRAMSOL].
+Models 2-4 were created by variably bending the first alphafold model manually at base-pair 34.
 (B) Pair-distance probability distribution functions $P(r)$ for these 4 conformations computed with CRYSOL [SVERGUN995].
@@
 The ‘data compare’ function in PRIMUS was then used to assess the statistical similarity of the datasets in reciprocal space.
+To focus on the $q$-range that gives rise to the features that we have observed in $P_{\mathrm{exp}}(r)$, the maximum value of $q$ of the data used for the test was set to 0.180.18\;Å-1.
+It computes a $p$-value by checking for patterns in the cross-correlation matrix between datasets, which does not depend on the values of the experimental errors [FRANKE2015].
+While $p>0.95$ means that two datasets are almost identical, low values of $p$, typically below a threshold of 0.010.050.01-0.05, indicate that the datasets are statistically independent.
+Such values are generally used to identify outliers, such as sample misloading, denaturing or radiation damage [MANALASTAS].
+Intermediate values of $p$ indicate some correlations between the datasets.
+Supplementary Table 1 shows that the $p$-values obtained by comparing the GAGE6 data successively with data recorded for GAGE6_1, GAGE6_2, GAGE6_3 are 0.493, 0.280, 0.078.
+This indicates decreasing correlations between the data sets, which are consistent with the nature of our samples.
+They derive from the same GAGE6 sequence, with only local sequence changes, growing from GAGE6_1 to GAGE6_3.
+Therefore one should not expect that such samples lead to completely statistically independent datasets as it occurs for outliers due to experimental problems.
+These $p$ values are interesting because our analysis (see Results section 3) detects a large-bending domain in GAGE6_1 and GAGE6_2 in a portion of the sequence which is not far from the bent domain of detected for GAGE6, which is compatible with a fairly large $p$-value whereas this strongly bent domain disappears for GAGE6_3, which should decrease the correlation with the GAGE6 dataset leading to a lower $p$-value.
+The statistical analysis appears therefore to provide a very good preview of the results presented in Sec.
+3.
 2.4 Derivation of $P_{\mathrm{exp}}(r)$ from experimental data
-The derivation of the pair-distance distribution function $P_{\mathrm{exp}}(r)$ from $I(q)$ versus $q$ via an indirect Fourier transform is performed with the program GNOM of the ATSAS package [MANALASTAS].
+The derivation of the pair-distance distribution function (Supplementary figure S3) $P_{\mathrm{exp}}(r)$ from $I(q)$ versus $q$ via an indirect Fourier transform is performed with the program GNOM of the ATSAS package [MANALASTAS].
 The program requires several input parameters which can variably affect the results.
@@
 As shown by the supplementary figure S5, varying $D_{\mathrm{max}}$ over a range of 2020\;Å (which is very significant in comparison with the length of the DNA sequences) only has a minute effect on the shape of $P_{\mathrm{exp}}(r)$.
+Furthermore, as small errors in $D_{\mathrm{max}}$ can alter $P(r)$ slightly [GRANT2015] persistence of features in the respective functions in the face of varying $D_{\mathrm{max}}$ serves as an additional indicator of the robustness of the solution.
 The choice of the regularisation parameter $\alpha$ is more delicate.
@@
 They are plotted in Fig.
-S5 of the supplementary material which also shows how the TQE of GNOM-4 depends on $\alpha$ for all the samples.
+S6 of the supplementary material which also shows how the TQE of GNOM-4 depends on $\alpha$ for all the samples.
 For GAGE6 the results of GNOM-5 and GNOM-4 are very close to each other for the shape of $P_{\mathrm{exp}}(r)$ and its estimated error bars.
@@
 Therefore, for some samples, the calculation of $P_{\mathrm{exp}}(r)$ raises some doubts concerning the validity of the results.
-However, supplementary figures S5 and S6 show that $\alpha$ essentially determines the magnitude of the humps of $P_{\mathrm{exp}}(r)$ but has little effect on their position with respect to $r$.
+However, Supplementary figures S6 and S7 show that $\alpha$ essentially determines the magnitude of the humps of $P_{\mathrm{exp}}(r)$ but has little effect on their position with respect to $r$.
 This is important because Fig.
@@
 Supplementary figure S10 shows that the same is true for all sequences.
-Simulated $P(r$) functions (Figure 2) were generated using CRYSOL (Svergun et al., 1995) on atomistic models, applying fake 1% errors to the theoretical intensity and proceeding with the standard derivation of $P(r)$ in Primus.
+This shows that SAXS data for flexible polymers, which can hardly be distinguished from each other in reciprocal space, even with mathematical tools, can contain detailed information which is revealed when the results are expressed in real space.
+The above factors together with the simulated $P(r)$ functions which show bending a dsDNA model will cause bumps to appear in the function, the fact that AT tracts have been proven to cause bending of dsDNA, the good fits of each $P(r)$ function to the data (Supplementary Figure S8) and their reasonable TQE (Total Quality Estimate) scores strongly support the validity of the information collected in real space.
+This is further confirmed by the correlation between the sequence of our DNA samples and the local properties that we detect in our analysis based on $P_{\mathrm{exp}}(r)$.
+Simulated $P(r$) functions (Figure 2) were generated using CRYSOL (Svergun et al., 1995) on the first Alphafold 3 model and the 3 subsequent manually bent models, applying fake 1% errors to the theoretical intensity and proceeding with the standard derivation of $P(r)$ in Primus.
 2.5 Analysis: from $P_{\mathrm{exp}}(r)$ to the conformation of the samples.
@@
 For this simple model, the theoretical calculation of the scattering function is possible within some limits and its comparison with the experimental data determines the appropriate model parameter, from which the persistence length can be derived.
-The same idea was refined by Schellmann [SCHELLMAN1974] for a better description of DNA flexibility which took into account dihedral angles $\phi$ between the planes defined by pairs of consecutive segments.
-However, to allow for analytical calculations, he only considered two limiting cases, either $\phi$ is constrained to zero, i.e.
-the model evolves in a plane, or $\phi$ is totally random.
-For DNA the reality lies between these two limits which correspond to either an infinite torsional rigidity ($\phi=0$) or zero torsional rigidity ($\phi$ random, evolving freely without any energy cost).
+The same idea was refined by Schellmann [SCHELLMAN1974] for a better description of DNA flexibility which took into account dihedral angles $\varphi$ between the planes defined by pairs of consecutive segments.
+However, to allow for analytical calculations, he only considered two limiting cases, either $\varphi$ is constrained to zero, i.e.
+the model evolves in a plane, or $\varphi$ is totally random.
+For DNA the reality lies between these two limits which correspond to either an infinite torsional rigidity ($\varphi=0$) or zero torsional rigidity ($\varphi$ random, evolving freely without any energy cost).
 This is why we have chosen an extended Kratky-Porod model with an added energy contribution for the dihedral angles in order to describe not only the persistence length of DNA but also its proper torsional rigidity.
+Figure 3: The DNA polymer model used in the analysis of the SAXS data.
 2.5.2 The DNA polymer model
@@
 3.
-It is not concerned with the internal structure of the DNA helix but only with the conformation of its backbone.
+It is not concerned with the internal structure of the DNA helix but only with the conformation of the line going through the center of the double helix.
 It consists of $N$ objects representing the base pairs, each separated by the base pair distance in DNA a=3.34a=3.34\;Å.
 Its N1N-1 segments can be viewed as a generalised Kratky-Porod model.
-Its conformation is defined by the bending angles $\theta_{n}$, $n=2,\ldots,N-1$ at sites $n$, between segment n1n-1 and $n$ and the dihedral angles $\phi_{n}$, $n=2,\ldots,N-2$, between the planes defined by segments [n1,n][n-1,n] and segments $[n,n+1]$.
-The thermal fluctuations of the backbone are controlled by the variation of its bending and torsional energies $E_{\theta}$ and $E_{\phi}$
-$H=E_{\theta}+E_{\phi}=\sum_{n=2}^{N-1}K(1-\cos\theta_{n})+\sum_{n=2}^{N-2}C(1-\cos\phi_{n})\;.$ (1)
+Its conformation is defined by the bending angles $\theta_{n}$, $n=2,\ldots,N-1$ at sites $n$, between segment n1n-1 and $n$ ($\pi<\theta_{n}\leq\pi$) and the dihedral angles $\varphi_{n}$, $n=2,\ldots,N-2$, between the planes defined by segments [n1,n][n-1,n] and segments $[n,n+1]$ ($\pi/2<\varphi_{n}\leq\pi/2$).
+The reference state for this model is a straight DNA helix, with 11 base pairs per turn, i.e.
+an angle of about $33^{\circ}$ between them.
+The $\theta$ and $\varphi$ angles describe the deviations with respect to this structure.
+The thermal fluctuations of the backbone are controlled by the variation of its bending and torsional energies $E_{\theta}$ and $E_{\varphi}$
+$H=E_{\theta}+E_{\varphi}=\sum_{n=2}^{N-1}K(1-\cos\theta_{n})+\sum_{n=2}^{N-2}C(1-\cos\varphi_{n})\;.$ (1)
 The constants $K$ and $C$ set the scale of the bending and torsional energies.
-They are assumed to be the same along the full model, and the bending and twist energies are minimal for $\theta=0$ and $\phi=0$, i.e.
+They are assumed to be the same along the full model, and the bending and twist energies are minimal for $\theta=0$ and $\varphi=0$, i.e.
 the model is homogeneous, with no permanent bending or torsion, in order to avoid any bias towards a specific conformation other that a straight, untwisted, DNA molecule.
-Figure 3: The DNA polymer model used in the analysis of the SAXS data.
 The numerical calculations use dimensionless variables.
@@
 The constant $C$ determines the torsional rigidity $\xi$ of the model in the linear limit, i.e.
-for small torsional angles $\phi$ for which $E_{\phi}\approx(1/2)C\phi^{2}$ and a torque $\tau=\partial E_{\phi}/\partial\phi=C\phi$.
-In this limit a DNA molecule of length $L$ behaves as an elastic rod whose torsional torque grows with torsion as $\tau=(\xi/L)\phi$.
+for small torsional angles $\varphi$ for which $E_{\varphi}\approx(1/2)C\varphi^{2}$ and a torque $\tau=\partial E_{\varphi}/\partial\varphi=C\varphi$.
+In this limit a DNA molecule of length $L$ behaves as an elastic rod whose torsional torque grows with torsion as $\tau=(\xi/L)\varphi$.
 For DNA measurements of $\xi$ vary from about $200\;\mathrm{pN\cdot nm^{2}}$ to $480\;\mathrm{pN\cdot nm^{2}}$.
@@
 Conformations of this model at room temperature are generated using Monte-Carlo simulations.
-The first segment $\mathbf{R_{1}}$ is chosen arbitrarily along the $z$ axis of our reference frame, and a second segment $\mathbf{R_{2}}$ is added in the $x,z$ plane by selecting a random value of $\theta_{2}$ in the range $\pi,\pi]$.
+The first segment $\mathbf{R_{1}}$ is chosen arbitrarily along the $z$ axis of our reference frame, and a second segment $\mathbf{R_{2}}$ is added in the $x,z$ plane by selecting a random value of $\theta_{2}$ in the range $\pi<\theta_{2}\leq\pi$.
 From two vectors we can generate a local frame $x^{\prime},y^{\prime},z^{\prime}$, using $\mathbf{R_{2}}$ as the $z^{\prime}$ axis, choosing an orthogonal $x^{\prime}$ axis in the $\mathbf{R_{1}}$$\mathbf{R_{2}}$ plane and completing by a third axis $y^{\prime}$ to get an orthonormal frame.
-A new segment is added in this local frame by selecting $\theta$ in the range $\pi,\pi]$ and the dihedral angle $\phi$ in the range $\pi/2,\pi/2]$, and so on until the $59$ segments corresponding to our $60$-base-pair DNA samples have been added.
-Each time a new segment has been defined the contribution $E_{\theta}+E_{\phi}$ that it adds to the energy is compared to $k_{B}T$ ($k_{B}T=1$ in our dimensionless units) and accepted or rejected with the Metropolis algorithm [METROPOLIS1953].
+A new segment is added in this local frame by selecting $\theta$ in the range $\pi<\theta\leq\pi$ and the dihedral angle $\varphi$ in the range $\pi/2<\varphi\leq\pi/2$, and so on until the $59$ segments corresponding to our $60$-base-pair DNA samples have been added.
+Each time a new segment has been defined the contribution $E_{\theta}+E_{\varphi}$ that it adds to the energy is compared to $k_{B}T$ ($k_{B}T=1$ in our dimensionless units) and accepted or rejected with the Metropolis algorithm [METROPOLIS1953].
 Once a conformation of $59$ segments has been completed the probability distribution $P(r)$ of the pair distances between its 60 beads is obtained by a calculation which directly derives from its definition.
@@
 We also show the average of $|\theta_{n}(i)|$ restricted to bending angles which exceed a threshold $\theta_{0}=5^{\circ}$ to stress the domains which are prone to large bending.
+The average diehdral angle $\overline{|\varphi_{n}(i)|}$ and the standard deviation $\sigma_{\varphi}$ are computed with equations similar to Eqs.
+(2) and (3).
 Figure 4 shows a typical pattern from SAXS results: it is symmetrical with respect to the centre of the sequence.
@@
 The two peaks may be real, or one of them may be an artefact because the statistical analysis mixes sequences with $n=p$ and others with n=60p+1n=60-p+1 so that we sum up data with a single peak on the right and data with a single peak on the left.
-Figure 4: Statistics of the local bending angles $|\theta_{n}(i)|$ over the 1000 best saved conformations for the GAGE6 sample: the red full line shows $\overline{|\theta_{n}(i)|}$.
-The black line with circles shows the same average, limited to the bending angles which are above $\theta_{0}=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$
-Table 1: Small Angle X-ray scattering data collection parameters
+Table 1: Sample information and characterization - Small Angle X-Ray Scattering parameters
 (A) Sample details GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
@@
 In beam sample cell Co-flow Co-flow Co-flow Co-flow
-Size Exclusion Chromatograpy SEC-SAS
+Size Exclusion Chromatograpy SEC-SAXS
 […]
@@
 SEC flow rate, mL/min 0.4 0.4 0.4 0.4
-(B) SAS Data Collection GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
+(B) SAXS Data Collection GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
 Data acquisition/reduction software Scatterbrain 2.82 Scatterbrain 2.82 Scatterbrain 2.82 Scatterbrain 2.82
@@
 frames 183194183-194 selected (12 1 second frames) frames 156162156-162 selected (7 1 second frames) frames 163164163-164 selected (2 1 second frames) frames 166172166-172 selected (7 1 second frames)
-(C) SAS-derived structural parameters GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
+(C) SAXS-derived structural parameters GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
 Methods/Software ATSAS 4.0 ATSAS 4.0 ATSAS 4.0 ATSAS 4.0
@@
 From chemical composition 37.101 37.105 37.111 37.116
-From SAS concentration independent method (volume of correlation method) 30.06 31.61 35.53 34.95
+From SAXS concentration independent method (volume of correlation method) 30.06 31.61 35.53 34.95
 (D) Scattering particle size GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
 (E) Data deposition
-SASBDB ID: SASDT72 SASDV87 SASDV97 SASDVA7
+SASBDB ID: SASDXH7 SASDV87 SASDV97 SASDVA7
 However, if the sequence has some intrinsic asymmetry this gives an additional piece of information which can be combined with the SAXS data to solve the dilemma by examining the conformations one by one and reversing some of them so that the peak of large $|\theta_{n}|$ is always on the same side, for instance to the right that we call “forward” conformation.
@@
 4.
-The mirror positions are $n^{\prime}_{1}=N-n_{2}+1$ and $n^{\prime}_{2}=N-n_{1}+1$.
+The mirror positions are $n^{\prime}_{1}=N-n_{2}+1$ and $n^{\prime}_{2}=N-n_{1}+1$ i.e.
+$n^{\prime}_{1}=13$ and $n^{\prime}_{2}=25$ for the example above.
 Then we scan again the 1000 best conformations and compute for each of them $s=\sum|\theta_{n}|,n=n_{1}\ldots n_{2}$ and $s^{\prime}=\sum|\theta_{n}|,n=n^{\prime}_{1}\ldots n^{\prime}_{2}$.
-If the difference between $s$ and $s^{\prime}$ is significant, then one of the two peaks is ficticious.
+If the difference between $s$ and $s^{\prime}$ is significant, which is evaluated by calculating $\Delta s/s=|s^{\prime}-s|/s$, then one of the two peaks is ficticious.
 If $s>s^{\prime}$ this conformation is already in the forward orientation.
@@
 As shown in the Results section, this is what happens for the 4 samples that we studied.
-When reversing a conformation, we also reverse the indices for the dihedral angles $\phi_{n}$ paying attention to the fact that, while $n$ varies from $2$ to N1N-1 for $\theta_{n}$, it varies from $2$ to N2N-2 for $\phi_{n}$ so that, instead of $n^{\prime}$ one has to use $n^{\prime\prime}=N-n$ for $\phi$.
+When reversing a conformation, we also reverse the indices for the dihedral angles $\varphi_{n}$ paying attention to the fact that, while $n$ varies from $2$ to N1N-1 for $\theta_{n}$, it varies from $2$ to N2N-2 for $\varphi_{n}$ so that, instead of $n^{\prime}$ one has to use $n^{\prime\prime}=N-n$ for $\varphi$.
 Besides the elimination of spurious signals, the orientation process improves the quality of the averaging because it becomes real averaging over the 1000 best configurations instead of the superposition of two sets of data averaged over about 500 conformations.
+Figure 4: Statistics of the local bending angles $|\theta_{n}(i)|$ over the 1000 best saved conformations for the GAGE6 sample: the red full line shows $\overline{|\theta_{n}(i)|}$.
+The black line with circles shows the same average, limited to the bending angles which are above $\theta_{0}=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$.
+The full brown vertical lines show the positions $n_{1}=36$, $n_{2}=48$ which have been used for the ’orientation’ of the conformations and the dashed brown vertical lines show the mirror positions $n^{\prime}_{1}=13$ and $n^{\prime}_{2}=25$.
 After this process has been completed we know that all the conformations generated by the model have the same orientation with respect to the sequence, i.e.
@@
 The analysis can then be extended to detect other properties of the sample.
-In the Results section, we show how the dihedral angles vary along the sequence, and we also make histograms $H(n,\theta)$ showing, for each site, the distribution of the bending angles among the 1000 best conformations, and similarly for $H(n,\phi)$.
+In the Results section, we show how the dihedral angles vary along the sequence, and we also calculate the probability $P(n,|\theta|)$ showing, for each site, the distribution of the bending angles among the 1000 best conformations, and similarly for $P(n,|\varphi|)$.
 These figures are very helpful to get a view of the likely conformations of the samples in solution and to try to understand why some sequences may be favourable for protein binding.
@@
 3 Results and Discussion
-The CHROMIXS chromatograms indicated the presence of one single main peak eluting in a similar position for all the duplexes, with some small shoulders on either side of the peak (Supplementary Figure S1).
+3.1 Small Angle X-ray Scattering data
+The chromatograms computed with the program CHROMIXS [CHROMIXS] indicated the presence of one single main peak eluting in a similar position for all the duplexes, with some small shoulders on either side of the peak (Supplementary figure S1).
 All eluted peaks had 260:280nm UV absorbance ratios consistent with that of pure dsDNA (Supplementary figure S1).
@@
 In the case of GAGE6_3, the variation in the predicted $R_{g}$ across the chosen frames was minimal, ranging from 49.350.649.3-50.6\,Å with an average of 50.150.1\,Å.
+The calculation of the molecular weight of the frames across the chromatography peak worked well for GAGE6_1, GAGE6_2, and GAGE6_3; however, it was unsuccessful for GAGE6, perhaps due to the fact that this experiment is slightly noisier than the others.
+As shown in Supplementary figure S2, the molecular weight of the frames across the peaks are in agreement with the expected molecular weight of a dsDNA duplex of this size ($\approx 37$kDa).
+This demonstrates that at least in the case of 3 of the 4 chromatograms, the samples are clearly monodisperse.
+Despite this calculation not working for the GAGE6 sample, the $R_{g}$ values of the chosen frames for the GAGE6 experiment are in close agreement, demonstrating that this sample is also likely to be monodisperse.
 All scattering shown as $\log_{10}(I)$ vs.
@@
 The details concerning SAXS parameters are summarised in Table 1.
-3.1 Small Angle X-ray Scattering data
 Figure 5: Scattering data and Guinier fits of the various GAGE6 oligonucleotides: $I(q)$ versus $q$ plot, Guinier plot ($\ln[I(q)]$ vs $q^{2}$), and Guinier residual plots for (A) GAGE6, (B) GAGE6_1, (C) GAGE6_2, (D) GAGE6_3.
@@
 The curve deduced from the selected conformations of the polymer model (Figure 6A) provides a very good description of the experimental curve in the whole domain where the polymer model is valid, indicating in this instance that the polymer model ensemble can very accurately reproduce $P_{\mathrm{exp}}(r)$.
-Figure 6: Sequence GAGE6: Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve).
+Figure 6: Sequence GAGE6: (A) Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve).
 The grey region, for r<25r<25\,Å is the low-$r$ domain in which the model cannot describe the internal structure of the double helix.
-The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6: Statistics of the local bending angles $\theta_{n}(i)$ and dihedral angles $\phi_{n}(i)$ over the 1000 best-saved conformations for the GAGE6 sample, after orientation of the conformations as discussed in the Materials and Methods section.
-The red full line shows the average over $i$ of $|\theta_{n}(i)|$.
-The black line with circles shows the same average, limited to the bending angles which are above $\theta=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$.
-The thin black line, without symbols, shows the average over $i$ of the dihedral angles and the thin black dash-dot line shows their standard deviation.(C) Sequence GAGE6: Three-dimensional plot of the normalised probability distribution function $H(n,\theta)$.
+The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6: Upper part: Statistics of the absolute value of the local bending angles $|\theta_{n}(i)|$ over the 1000 best-saved conformations for the GAGE6 sample, after orientation of the conformations as discussed in the Materials and Methods section.
+The red full line shows the average over $i$ of $|\theta_{n}(i)|$ versus $n$.
+The black line with circles shows the same average, limited to the bending angles which are above $|\theta|=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$.
+Lower part: Statistics of the absolute values of the dihedral angles The red full line shows the average over $i$ of $|\varphi_{n}(i)|$ versus $n$ and the the dashed blue curve shows their standard deviation.
+(C) Sequence GAGE6: Three-dimensional plot of the normalised probability distribution function $P(n,|\theta|)$.
 The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$.
-The value $W$ marked on the figure shows the fraction of conformations which have a bending angle above the value shown below $W$ within the domain indicated by the two values of $n$.
-Figure 6B shows the statistics of the bending angles $|\theta_{n}|$ and dihedral angles $\phi_{n}$ with their averages over conformations and standard deviations.
+The value $W$ marked on the figure shows the fraction of bending angles $|\theta_{n}|$ above the value shown below $W$ within the domain indicated by the two values of $n$.
+Figure 6B shows the statistics of the bending angles $|\theta_{n}|$ and dihedral angles $|\varphi_{n}|$ with their averages over conformations and theit standard deviations.
 For the bending angles this figure is analogous to Fig.
@@
 Their differences were significant, suggesting that one of the two mirror images was fictitious, which is confirmed by the figure as one of them has been fully erased in the orientation step.
-The three-dimensional plot of Fig. 6C shows the probability distribution of the bending angles $|\theta|$ for each site $n$ along the sequence. These distributions are normalised
-$\int_{0}^{\pi}H(n,\theta)=1\quad{\mathrm{for~all~}}n$ (5)
-which allows us to calculate the fraction $W$ of conformations which have a bending angle above some value in a given domain.
+The three-dimensional plot of Fig. 6C shows the probability $P(n,|\theta|)$ of the bending angles $|\theta_{n}|$ for each site $n$ along the sequence. These distributions are normalised
+$\int_{0}^{\pi}P(n,|\theta|)\;\mathrm{d}|\theta|=1\quad{\mathrm{for~each~}}n$ (5)
+which allows us to calculate the fraction $W$ of the bending angles $|\theta_{n}|$ above a some value in a given domain.
 The result $W=0.06557$ is shown in Fig 6C for $35\leq n\leq 49$ and a lower value of $\theta$ of $45^{\circ}$.
-Thus, although most of the bending angles are well below $30^{\circ}$, in some regions of the sequence there are about $6\%$ of the conformations which show bending above $45^{\circ}$.
+Thus, although most of the bending angles are well below $30^{\circ}$, in some regions of the sequence there are about $6\%$ of the bending angles above $45^{\circ}$.
 The statistics clearly show that one region of the sequence is special, with a much higher flexibility and larger bending than elsewhere.
-The colour view of the three-dimensional histogram Fig.
-6C, (see Supplementary material Fig.
-S8) suggests that the large bending angles for $35\leq n\leq 49$ are separated from the bulk of smaller bending angles, as if they result from some permanent bend in this region and not simply from large amplitude thermal fluctuations.
-As this part of the histogram only concerns $6\%$ of the conformations, the resolution of the analysis of the SAXS data does not however allow us to make a definitive statement on this point.
-We would like to stress that these statistical results which provide precise information on the local properties of the sample, are extracted from a probability distribution function which looks rather featureless although it results from a large ensemble of pair distances.
+The colour view of the three-dimensional probability $P(n,|\theta|)$ (Fig.
+6C), (see Supplementary material Fig.
+S9) suggests that the large bending angles for $35\leq n\leq 49$ are separated from the bulk of smaller bending angles, as if they result from some permanent bend in this region and not simply from large amplitude thermal fluctuations.
+As this part of the probability distribution only concerns $6\%$ of the bending angles, the resolution of the analysis of the SAXS data does not however allow us to make a definitive statement on this point.
+We would like to stress that these statistical results which provide precise information on the local properties of the sample, are extracted from a probability distribution function $P_{\mathrm{exp}}(r)$ which looks rather featureless although it results from a large ensemble of pair distances.
 It is remarkable that this smooth looking, almost straight function, actually encodes so much information.
@@
 More recently Theodorakoupolous and Peyrard [DNAFLEX] reviewed this concept and theorised that below the melting point of a duplex, an increase in temperature towards physiological temperature may increase the opening of base pairs enough to act as flexible hinges in dsDNA, which may have wider implications for DNA structure and recognition by proteins.
-Figure 7: Sequence GAGE6_1: Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve).
+Figure 7: Sequence GAGE6_1: (A) Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve).
 The grey region, for r<25r<25\,Å is the low-$r$ domain in which the model cannot describe the internal structure of the double helix.
-The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_1: Statistics of the local bending angles $\theta_{n}(i)$ and dihedral angles $\phi_{n}(i)$ over the 1000 best-saved conformations for the GAGE6_1 sample, after orientation of the conformations as discussed in the Materials and Methods section.
-The red full line shows the average over $i$ of $|\theta_{n}(i)|$.
-The black line with circles shows the same average, limited to the bending angles which are above $\theta=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$.
-The thin black line, without symbols, shows the average over $i$ of the dihedral angles and the thin black dash-dot line shows their standard deviation.(C) Sequence GAGE6_1: Three-dimensional plot of the normalised probability distribution function $H(n,\theta)$.
+The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_1: Upper part: Statistics of the absolute value of the local bending angles $|\theta_{n}(i)|$ over the 1000 best-saved conformations for the GAGE6 sample, after orientation of the conformations as discussed in the Materials and Methods section.
+The red full line shows the average over $i$ of $|\theta_{n}(i)|$ versus $n$.
+The black line with circles shows the same average, limited to the bending angles which are above $|\theta|=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$.
+Lower part: Statistics of the absolute values of the dihedral angles The red full line shows the average over $i$ of $|\varphi_{n}(i)|$ versus $n$ and the the dashed blue curve shows their standard deviation.
+(C) Sequence GAGE6_1: Three-dimensional plot of the normalised probability distribution function $P(n,|\theta|)$.
 The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$.
-The value $W$ marked on the figure shows the fraction of conformations which have a bending angle above the value shown below $W$ within the domain indicated by the two values of $n$.
+The value $W$ marked on the figure shows the fraction of bending angles $|\theta_{n}|$ above the value shown below $W$ within the domain indicated by the two values of $n$.
 Alongside the fluctuational opening of the duplex, further unwinding of DNA regions is also important.
@@
 The dihedral angles show the same behaviour, with large twist fluctuations being restricted to the same narrow domain.
-Figure 7C confirms the presence of large bending angles in a narrow domain and the integration of $H(n,\theta)$ for $41\leq n\leq 52$ and $\theta>45^{\circ}$ shows that the fraction $W$ of bending angles above $45^{\circ}$ has drastically increased from 6.4% to 12.6%.
+Figure 7C confirms the presence of large bending angles in a narrow domain and the integration of $P(n,|\theta|)$ for $41\leq n\leq 52$ and $\theta>45^{\circ}$ shows that the fraction $W$ of bending angles $|\theta_{n}|$ above $45^{\circ}$ has drastically increased from 6.4% to 12.6%.
 The GAGE6_1 sequence therefore has a high probability of being sharply bent in a narrow region.
@@
 This strengthens the notion that $P(r)$ has a high sensitivity to molecular conformations, particularly bending, with sharp bending generating a function with two bumps.
-Figure 8: Sequence GAGE6_2: Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve).
+Figure 8: Sequence GAGE6_2: (A) Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve).
 The grey region, for r<25r<25\,Å is the low-$r$ domain in which the model cannot describe the internal structure of the double helix.
-The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_2: Statistics of the local bending angles $\theta_{n}(i)$ and dihedral angles $\phi_{n}(i)$ over the 1000 best-saved conformations for the GAGE6_2 sample, after orientation of the conformations as discussed in the Materials and Methods section.
-The red full line shows the average over $i$ of $|\theta_{n}(i)|$.
-The black line with circles shows the same average, limited to the bending angles which are above $\theta=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$.
-The thin black line, without symbols, shows the average over $i$ of the dihedral angles and the thin black dash-dot line shows their standard deviation.(C) Sequence GAGE6_2: Three-dimensional plot of the normalised probability distribution function $H(n,\theta)$.
+The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_2: Upper part: Statistics of the absolute value of the local bending angles $|\theta_{n}(i)|$ over the 1000 best-saved conformations for the GAGE6 sample, after orientation of the conformations as discussed in the Materials and Methods section.
+The red full line shows the average over $i$ of $|\theta_{n}(i)|$ versus $n$.
+The black line with circles shows the same average, limited to the bending angles which are above $|\theta|=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$.
+Lower part: Statistics of the absolute values of the dihedral angles The red full line shows the average over $i$ of $|\varphi_{n}(i)|$ versus $n$ and the the dashed blue curve shows their standard deviation.
+(C) Sequence GAGE6_2: Three-dimensional plot of the normalised probability distribution function $P(n,|\theta|)$.
 The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$.
-The value $W$ marked on the figure shows the fraction of conformations which have a bending angle above the value shown below $W$ within the domain indicated by the two values of $n$.
+The value $W$ marked on the figure shows the fraction of bending angles $|\theta_{n}|$ above the value shown below $W$ within the domain indicated by the two values of $n$.
 In GAGE6_2, in addition to replacing the AT base pairs between positions $28\leq n\leq 31$ with GC pairs, the AT pairs in positions $38\leq n\leq 44$ have now also been replaced with GC pairs.
@@
 This points out again that, although sequence and flexibility are correlated, the preferred conformations of a DNA sequence in solution are not only determined by the local sequence but collective effects at larger distances are also important.
-The histograms of the bending angles shown in Fig.
+The probability distribution of the bending angles $P(n,|\theta|)$ shown in Fig.
 8C and Supplementary Fig.
-S8 are somewhat similar to the data from Fig.
-7: the fraction of large bending angles in the range $44\leq n\leq 57$ is of the order of 9.6%, suggesting that GAGE6_2 is also sharply bent in a narrow region.
-Figure 9: Top panel: sequence GAGE6_2.
-Comparison of AutoGNOM-4-$P_{\mathrm{exp}}(r)$ (small circle points and thin black curve with error bars) and $P(r)$ given by the 1000 best conformations of the polymer model (thick red curve).
-The gray region for r<25r<25\;Å is the low-r domain in which the model cannot describe the internal structure of the double helix.
-Bottom panel: Statistics of the local bending angles $\theta_{n}(i)$ over the 1000 best model conformations: the red full line shows the average over $i$ of $|\theta_{n}(i)|$.
-The thin black line without symbols shows the same average limited to the bending angles $\theta$ above $5^{\circ}$ and dashed blue curve shows the standard deviations of $\theta_{n}(i)$.
-The thin black line without symbols shows the average of the dihedral angles and the thin black-dot-line shows their standard deviation.
-The vertical dashed lines show the boundaries of the domain where large bending was observed in the analysis of AutoGNOM-5-$P_{\mathrm{exp}}(r)$.
+S9 are somewhat similar to the data from Fig.
+7: the fraction of large bending angles $|\theta_{n}|$ in the range $44\leq n\leq 57$ is of the order of 9.6%, suggesting that GAGE6_2 is also sharply bent in a narrow region.
 However, the results shown in Fig.
@@
 9.
+Figure 9: Top panel: sequence GAGE6_2.
+Comparison of AutoGNOM-4-$P_{\mathrm{exp}}(r)$ (small circle points and thin black curve with error bars) and $P(r)$ given by the 1000 best conformations of the polymer model (thick red curve).
+The gray region for r<25r<25\;Å is the low-r domain in which the model cannot describe the internal structure of the double helix.
+Bottom panel: Upper part: Statistics of the absolute value of the local bending angles $|\theta_{n}(i)|$ over the 1000 best-saved conformations for the GAGE6 sample, after orientation of the conformations as discussed in the Materials and Methods section.
+The red full line shows the average over $i$ of $|\theta_{n}(i)|$ versus $n$.
+The black line with circles shows the same average, limited to the bending angles which are above $|\theta|=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$.
+Lower part: Statistics of the absolute values of the dihedral angles The red full line shows the average over $i$ of $|\varphi_{n}(i)|$ versus $n$ and the the dashed blue curve shows their standard deviation.
+The vertical dashed lines show the boundaries of the domain where large bending was observed in the analysis of AutoGNOM-5-$P_{\mathrm{exp}}(r)$.
 The top panel of Fig.
@@
 Figure 10B shows a broad domain with large bending angles with a sharp maximum at position $n=26$ which coincides with two AT pairs in the middle of a GC sequence.
-The three-dimensional plot of the histograms $H(n,\theta)$ plotted in Fig.
+The three-dimensional plot of the probability $P(n,|\theta|)$ plotted in Fig.
 10C and Supplementary Fig.
-S8 shows larger bending angles in the domain $17\leq n\leq 35$ but the integrated weight of the histograms for angles exceeding $45^{\circ}$ in this region shows that only 3.8% of the conformations exhibit large bending.
+S9 shows larger bending angles in the domain $17\leq n\leq 35$ but the integrated weight of the probability $P(n,|\theta|)$ for angles exceeding $45^{\circ}$ in this region shows that only 3.8% of the angles correspond to large bending.
 As suggested by its sequence, mostly GC base pairs and only a few small domains with AT pairs, not wider than three consecutive sites, the GAGE6_3 sequence has mostly straight conformations.
-Figure 10: Sequence GAGE6_3: Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve).
+Figure 10: Sequence GAGE6_3: (A) Comparison between $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (small circle points and thin black curve with error bars) and $P(r)$ calculated for the 1000 best conformations generated theoretically (thick red curve).
 The grey region, for r<25r<25\,Å is the low-$r$ domain in which the model cannot describe the internal structure of the double helix.
-The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_3: Statistics of the local bending angles $\theta_{n}(i)$ and dihedral angles $\phi_{n}(i)$ over the 1000 best-saved conformations for the GAGE6_3 sample, after orientation of the conformations as discussed in the Materials and Methods section.
-The red full line shows the average over $i$ of $|\theta_{n}(i)|$.
-The black line with circles shows the same average, limited to the bending angles which are above $\theta=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$.
-The thin black line, without symbols, shows the average over $i$ of the dihedral angles and the thin black dash-dot line shows their standard deviation.(C) Sequence GAGE6_3: Three-dimensional plot of the normalised probability distribution function $H(n,\theta)$.
+The thick black dashed line shows the function $P_{\mathrm{exp}}(r)$ obtained with AutoGNOM-4 (see section 2.4) (B) Sequence GAGE6_3: Upper part: Statistics of the absolute value of the local bending angles $|\theta_{n}(i)|$ over the 1000 best-saved conformations for the GAGE6 sample, after orientation of the conformations as discussed in the Materials and Methods section.
+The red full line shows the average over $i$ of $|\theta_{n}(i)|$ versus $n$.
+The black line with circles shows the same average, limited to the bending angles which are above $|\theta|=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$.
+Lower part: Statistics of the absolute values of the dihedral angles The red full line shows the average over $i$ of $|\varphi_{n}(i)|$ versus $n$ and the the dashed blue curve shows their standard deviation.
+(C) Sequence GAGE6_3: Three-dimensional plot of the normalised probability distribution function $P(n,|\theta|)$.
 The data are not shown for small bending angles, for which the probabilities are the highest, to better show the data at larger $\theta$.
-The value $W$ marked on the figure shows the fraction of conformations which have a bending angle above the value shown below $W$ within the domain indicated by the two values of $n$.
+The value $W$ marked on the figure shows the fraction of bending angles $|\theta_{n}|$ above the value shown below $W$ within the domain indicated by the two values of $n$.
 Supplementary figure S10 shows the results of the analysis of the functions $P_{\mathrm{exp}}(r)$ obtained with Auto-GNOM-4 for the four samples (see supplementary figure Fig.
-S5).
+S6).
 It can be compared with Figs.
@@
 Supplementary Materials:
-Figure 1: Supplementary Figure S1: Figure S1 shows the absorbance traces of all the duplexes at 214nm, 260nm and 280nm vs time (seconds) followed by scattering traces of all the duplexes plotted from CRHOMIXS showing average X-ray intensity against frame number.
+Supplementary Figure S1: Figure S1 shows the absorbance traces of all the duplexes at 214nm, 260nm and 280nm vs time (seconds) followed by scattering traces of all the duplexes plotted from CRHOMIXS showing average X-ray intensity against frame number.
 Predicted Rg values are shown as blue curves and were predicted using CHROMIXS.
@@
 GAGE6 - - - -
-GAGE6_1 0.614 - - -
-GAGE6_2 0.614 0.855 - -
-GAGE6_3 0.209 0.855 0.614 -
-Figure 2: Supplementary Figure S2: Comparative distance distribution functions of the different oligonucleotides used in this study.
+GAGE6_1 0.493 - - -
+GAGE6_2 0.280 0.493 - -
+GAGE6_3 0.078 0.493 0.753 -
+Supplementary Table 1: Correlation Map Test P-values after regridding the data truncated to $q\leq 0.18\;$Å-1 which is the domain that we use in the analysis.
+Supplementary Figure S2: (A) The GAGE6_1 mean intensity vs.
+frame number chromatogram (pink) shown alongside the results of the MoW calculation (black squares with light blue error bars) as performed by BioXTAS RAW.
+The expected mass of a 60bp dsDNA duplex is shown as a grey dotted horizontal line.
+The frames chosen for downstream analysis of P(r) are highlighted by a light green box.
+(B) The same as above but for GAGE6_2.
+(C) The same as above but for GAGE6_3.
+Supplementary Figure S3: Comparative distance distribution functions AutoGNOM-5-$P_{\mathrm{exp}}(r)$ of the different oligonucleotides used in this study.
 Sequence of each oligonucleotide indicated above $P_{\mathrm{exp}}(r)$ function with AT tracts highlighted in red and GC tracts in green.
 Sequence-dependent changes in the $P_{\mathrm{exp}}(r)$ function indicate the position of multiple bumps in the AT tract containing duplexes and a smooth descent for GAGE6_3.
-Figure 3: Supplementary Figure S3: Comparative $P(r)$ functions for datasets accompanied by 1x (pink line) and 2x (black line) standard error show consistency in the shape of P(r).
+Supplementary Figure S4: Comparative AutoGNOM-5-$P_{\mathrm{exp}}(r)$ functions for datasets accompanied by 1x (pink line) and 2x (black line) standard error show consistency in the shape of P(r).
 Errors bars only plotted for data from 1x standard error (pink).
-Figure 4: Supplementary Figure S4: Variation of $D_{\mathrm{max}}$ parameter around the reported values shows that $P(r)$ features remain consistently present.
-Figure 5: Supplementary Figure S5: Comparison between the functions $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (red curve with error bars) and with AutoGNOM-4 (black curve with error bars) for the four DNA samples.
+Supplementary Figure S5: Variation of $D_{\mathrm{max}}$ parameter around the reported values shows that AutoGNOM-5-$P_{\mathrm{exp}}(r)$ features remain consistently present.
+Supplementary Figure S6: Comparison between the functions $P_{\mathrm{exp}}(r)$ calculated with AutoGNOM-5 (red curve with error bars) and with AutoGNOM-4 (black curve with error bars) for the four DNA samples.
 The values of the regularisation parameter $\alpha$ used for each curve are indicated inside each panel.
 The inset panels show the variation of the Total Quality Estimate (TQE) versus $\alpha$ (in logarithmic scale) calculated by AutoGNOM-4.
-Figure 6: Supplementary Figure S6: Variation of the GNOM alpha parameter around the fitted value indicates that while the magnitude of features changes, their position relative to r remains consistent.
-Figure 7: Supplementary Figure S7: Plots of the GNOM fits to the scattering data and their corresponding residual plots.
-Figure 8: Supplementary Figure S8: Another view of the three-dimensional histograms $H(n,\theta)$ which highlights some features that are difficult to see on the three-dimensional plots.
-The left part shows the projection of the histograms for each site on a single plane which displays the probability of a given value of $\theta$, whatever the position in the sequence: $H_{\mathrm{sum}}(\theta)=\sum_{n=2}^{N-1}H(n,\theta)$ The right part uses a color code to show the three-dimensional histogram in a two-dimensional image.
-This is easier to get an idea of the map of the bending angles along the sequence.
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Figure 9: Supplementary figure S9: Analysis of the $P_{\mathrm{exp}}(r)$ obtained with Auto-GNOM-4 (see Fig.
-S5).
+Supplementary Figure S7: Variation of the GNOM alpha parameter around the fitted value indicates that while the magnitude of features changes, their position relative to r remains consistent.
+Supplementary Figure S8: Plots of the GNOM fits to the scattering data and their corresponding residual plots.
+Supplementary Figure S9: Another view of the three-dimensional probabilities $P(n,|\theta|)$ which highlights some features that are difficult to see on the three-dimensional plots.
+The left part shows the sum over $n$ of the projections of $P(n,|\theta|)$ on a single plane which displays the probability of a given value of $|\theta|$, whatever the position in the sequence:
+$\rho(\theta)=\sum_{n=2}^{N-1}P(n,|\theta|)$ (6)
+The right part uses a color code to show the three-dimensional probability in a two-dimensional image. This is easier to get an idea of the map of the bending angles along the sequence.
+[Uncaptioned image] [Uncaptioned image]
+[Uncaptioned image] [Uncaptioned image]
+Supplementary figure S10: Analysis of the $P_{\mathrm{exp}}(r)$ obtained with Auto-GNOM-4 (see Fig.
+S6).
 For each sample the top panel shows the function $P_{\mathrm{exp}}(r)$ with error bars (black) and its fit by the polymer model (red full line).